### v2

# On a fuzzy Landau Equation: Part I. A variational approach

Abstract
In this paper, we introduce a fuzzy Landau equation, where particles interact through delocalised Coulomb collisions. We establish a variational characterisation that recasts the fuzzy Landau equation within the framework of GENERIC systems (General Equations for Non-Equilibrium Reversible-Irreversible Coupling), obtaining an equivalent characterisation of $\mathcal{H}$-solutions.
1 Introduction
In this paper, we consider the inhomogeneous (kinetic) fuzzy Landau equation:
$\partial_{t}f+v\cdot\nabla_{x}f=Q_{\sf fuz}(f,f)\quad\text{on}\quad[0,T]\times\Omega\times\mathbb{R}^{d},~\text{and}~f(0,x,v)=f_{0}(x,v).$ (1.1)
The spatial domain $\Omega$ can be $\mathbb{T}^{d}$ or $\mathbb{R}^{d}$ ($d\geq 2$). This equation characterises the time evolution of the unknown $f_{t}(x,v):[0,T]\times\Omega\times\mathbb{R}^{d}\to\mathbb{R}_{+}$, describing the distribution of particles in a plasma at time $t$ at position $x$ and with velocity $v$. The initial datum $f_{0}$ is a given distribution on the phase space. The fuzzy Landau collision term is given by
$Q_{\sf fuz}(f,f)=$ $\nabla_{v}\cdot\Big(\int_{\Omega\times\mathbb{R}^{d}}\kappa(x-x_{*})a(v-v_{*})\big(f_{*}\nabla_{v}f-f\nabla_{v_{*}}f_{*}\big){\,\rm d}x_{*}{\,\rm d}v_{*}\Big),$
where $f_{*}=f(x_{*},v_{*})$ denotes the density of particle at different position and velocity $(x_{*},v_{*})$. The kernel $a:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ is given by
$a(z)=A(z)\Pi_{z^{\perp}},$ (1.2)
where $A:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is a weight, and $\Pi_{z^{\perp}}$ denotes the projection operator onto $z^{\perp}$,
$\Pi_{z^{\perp}}=\operatorname{Id}-\frac{z\otimes z}{|z|^{2}}.$
When $A(z)=|z|^{2+\gamma}$, $a$ is known as the Coulomb interaction kernel,
$a(z)=|z|^{2+\gamma}\Pi_{z^{\perp}}=|z|^{\gamma}(|z|^{2}\operatorname{Id}-z\otimes z).$ (1.3)
In (1.3), $\gamma$ is a physical parameter, where $\gamma>0$ corresponds to the so-called hard potentials and $\gamma<0$ to the soft potentials (with a further classification of $2<\gamma<0$ as the moderately soft potentials and $d<\gamma\leq-2$ as the very soft potentials). In particular, the $\gamma=0$ is known as the Maxwellian case while the case $\gamma=-d$ is known as the Coulomb case. In this paper, we consider
$\gamma\in(-\gamma_{d},1],$
where $\gamma_{d}=\min(d,4)$. The condition $\gamma>-d$ ensures the local integrability of $|z|^{\gamma}$, and the solutions to the fuzzy Landau equations are well-defined for $\gamma>-4$, as shown in Part II [DH25].
One relevant choice of the spatial kernel $\kappa:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is $\kappa=\kappa^{\sigma}$, where for a fixed parameter $\sigma\in(0,1)$, $\kappa^{\sigma}$ is a renormalised kernel proportional to $\exp(-\langle x/{\sqrt{\sigma}}\rangle)$, where $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese bracket. With this choice of the spatial kernel, the fuzzy Landau equation can be viewed as an intermediate model between the spatially homogeneous Landau equation and the classical inhomogeneous Landau equation. In fact, by taking the limit $\sigma\to+\infty$, one recovers the homogeneous Landau equation, see Remark 2.4 below. By taking the limit $\sigma\to 0$ instead, the spatial kernel $\kappa^{\sigma}$ converges to a Dirac measure. In this case, the fuzzy Landau equation (1.1) formally converges to the classical inhomogeneous Landau equation, which is given by
$\left\{\begin{aligned} &\partial_{t}f+v\cdot\nabla_{x}f=Q_{\sf class}(f,f)\\ &Q_{\sf class}(f,f)=\nabla_{v}\cdot\Big(\int_{\mathbb{R}^{d}}a\big(f(x,v_{*})\nabla_{v}f(x,v)-f(x,v)\nabla_{v_{*}}f(x,v_{*})\big){\,\rm d}v_{*}\Big).\end{aligned}\right$. (1.4)
The classical Landau equation, which is an important partial differential equation in kinetic theory, describes the evolution of a particle distribution in a plasma, accounting for both free particle transportation and Coulomb collisions. The classical Landau collisions in (1.4) are localised in the sense that two particles with velocities $v$ and $v_{*}\in\mathbb{R}^{d}$ collide at the same position $x\in\Omega$.
In the case of the fuzzy Landau equation (1.1), we allow delocalised collisions. This means that two particles can collide with different $(x,v)$ and $(x_{*},v_{*})$, where the positions are restricted by the spatial kernel $\kappa$. Various interaction kernels $A$ and spatial kernels $\kappa$ will be discussed. The precise scope can be found in Assumption 2.1 below.
In this article, which is the first part of our series, we show that the fuzzy Landau equation (1.1) can be rigorously cast into the so-called GENERIC framework via a variational formulation, obtaining an equivalent characterisation of $\mathcal{H}$-solutions.
The rest of this introduction is organised as follows: In Section 1.1, we recall the GENERIC framework. In Section 1.2, we review the related work to the fuzzy Landau equation (1.1). In Section 1.3, we present and discuss the main results.
1.1 GENERIC structure
The GENERIC (General Equations for Non-Equilibrium Reversible Irreversible Coupling) framework provides a systematic method to derive thermodynamically consistent evolution equations describing complex systems consisting of both Hamiltonian (reversible) dynamics and gradient (irreversible) flows. It was first introduced in the context of complex fluids [GÖ97, GÖ97a]. Over the last decade, it has proved to be a versatile modelling tool for a variety of complex systems in physics and engineering, ranging from anisotropic inelastic solids to viscoplastic solids and thermoelastic dissipative materials, to name just a few: We refer the reader to the monographs [Ött05, PKG18] for a detailed exposition of this framework. However, a mathematically rigorous foundation for GENERIC is still under-developed. The present work contributes towards this overarching goal.
A GENERIC system describes the evolution of an unknown $\mathsf{z}$ in a state space $\mathsf{Z}$ via the equation
$\partial_{t}\mathsf{z}=\mathsf{L}{\,\rm d}\mathsf{E}+\mathsf{M}{\,\rm d}\mathsf{S}.$ (1.5)
In this equation:
  • The functionals $\mathsf{E},\mathsf{S}:\mathsf{Z}\to\mathbb{R}$ are interpreted as, energy and entropy functionals respectively; and ${\,\rm d}\mathsf{E}$, ${\,\rm d}\mathsf{S}$ are their differentials.
  • $\mathsf{L}(\mathsf{z})$, $\mathsf{z}\in\mathsf{Z}$ are antisymmetric operators mapping cotangent to tangent vectors and satisfying the Jacobi identity
    $\{\{\mathsf{G}_{1},\mathsf{G}_{2}\}_{\mathsf{L}},\mathsf{G}_{3}\}_{\mathsf{L}}+\{\{\mathsf{G}_{2},\mathsf{G}_{3}\}_{\mathsf{L}},\mathsf{G}_{1}\}_{\mathsf{L}}+\{\{\mathsf{G}_{3},\mathsf{G}_{1}\}_{\mathsf{L}},\mathsf{G}_{2}\}_{\mathsf{L}}=0,$
    for all functions $\mathsf{G}_{i}:\mathsf{Z}\to\mathbb{R}$, $i=1,2,3$, where the Poisson bracket $\{\cdot,\cdot\}_{\mathsf{L}}$ is defined as follows:
    $\{\mathsf{F},\mathsf{G}\}_{\mathsf{L}}={\,\rm d}\mathsf{F}\cdot\mathsf{L}{\,\rm d}\mathsf{G}.$
    Here, $\cdot$ signifies the duality pairing of the cotangent and tangent spaces of $\mathsf{Z}$.
    The antisymmetry of $\mathsf{L}$ means that $\mathsf{v}\cdot\mathsf{L}\mathsf{w}=-\mathsf{w}^{*}\cdot\mathsf{L}\mathsf{v}^{*}$, where $\mathsf{v}^{*},\mathsf{w}^{*}$ refer to the dual vectors of $\mathsf{v},\mathsf{w}$;
  • $\mathsf{M}(\mathsf{z})$, $\mathsf{z}\in\mathsf{Z}$ is a symmetric and positive semi-definite operators mapping cotangent to tangent vectors in the sense that
    $\mathsf{v}\cdot\mathsf{M}\mathsf{w}=\mathsf{w}^{*}\cdot\mathsf{M}\mathsf{v}^{*},\quad\mathsf{v}^{*}\cdot\mathsf{M}\mathsf{v}\geq 0\quad\text{for any }\mathsf{v},\,\mathsf{w};$
  • The following degeneracy conditions are satisfied:
    $\mathsf{L}(\mathsf{z}){\,\rm d}\mathsf{S}(\mathsf{z})=0\quad\text{and}\quad\mathsf{M}(\mathsf{z}){\,\rm d}\mathsf{E}(\mathsf{z})=0\quad\text{for all }\mathsf{z}.$ (1.6)
Note that pure Hamiltonian systems and pure (dissipative) gradient flow systems are special cases of GENERIC corresponding to $\mathsf{M}\equiv 0$ and $\mathsf{L}\equiv 0$, respectively.
The conditions satisfied by the building blocks $\{\mathsf{E},\mathsf{S},\mathsf{L},\mathsf{M}\}$ ensure that in any solution to (1.5), the energy $\mathsf{E}$ is conserved and the entropy $\mathsf{S}$ is non-decreasing. In fact,
$\frac{{\,\rm d}}{{\,\rm d}t}\mathsf{E}(\mathsf{z}_{t})=\partial_{t}z\cdot{\,\rm d}\mathsf{E}=(\mathsf{L}{\,\rm d}\mathsf{E}+\mathsf{M}{\,\rm d}\mathsf{S})\cdot{\,\rm d}\mathsf{E}=\underbrace{\mathsf{L}{\,\rm d}\mathsf{E}\cdot{\,\rm d}\mathsf{E}}_{=0}+\underbrace{\mathsf{M}{\,\rm d}\mathsf{S}\cdot{\,\rm d}\mathsf{E}}_{=0}=0,$
and by a similar computations,
$\frac{{\,\rm d}}{{\,\rm d}t}\mathsf{S}(\mathsf{z}_{t})={\,\rm d}\mathsf{S}\cdot\mathsf{M}{\,\rm d}\mathsf{S}\geq 0.$
Thus, the first and second laws of thermodynamics are automatically satisfied for GENERIC systems. This, together with the structures of energy and entropy functionals strongly suggest a connection between the GENERIC framework and variational principles. However, while this idea was rigorously established for pure Hamiltonian systems (see for instance, [Arn13]) or pure gradient-flow systems, see for instance [AGS08], it is still underdeveloped for full GENERIC systems. This is one of the key challenges in the field, see [Ött18a].
The first author, Peletier and Zimmer [DPZ13] propose a variational characterisation: a curve $\mathsf{z}:[0,T]\to\mathsf{Z}$ is a solution to (1.5) if and only if $J(\mathsf{z})=0$, where
$J(\mathsf{z})=\mathsf{S}(\mathsf{z}_{0})-\mathsf{S}(\mathsf{z}_{T})+\frac{1}{2}\int_{0}^{T}\|\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E}\|_{\mathsf{M}^{-1}}^{2}{\,\rm d}t++\frac{1}{2}\int_{0}^{T}\|{\,\rm d}\mathsf{S}\|_{\mathsf{M}}^{2}{\,\rm d}t,$ (1.7)
and the (formal) inner products associated with $\mathsf{M}$ and $\mathsf{M}^{-1}$ are as follows:
$(a,b)_{\mathsf{M}}=a\cdot\mathsf{M}b\quad\text{and}\quad(a,b)_{\mathsf{M}^{-1}}=a\cdot\mathsf{M}^{-1}b.$
This variational characterisation is motivated by the following computations:
$J(\mathsf{z})$ $=\mathsf{S}(\mathsf{z}_{0})-\mathsf{S}(\mathsf{z}_{T})+\frac{1}{2}\int_{0}^{T}\big(\|\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E}\|_{\mathsf{M}^{-1}}^{2}+\|{\,\rm d}\mathsf{S}\|_{\mathsf{M}}^{2}\big){\,\rm d}t$
$=-\int_{0}^{T}(\partial_{t}\mathsf{z},\mathsf{M}{\,\rm d}\mathsf{S})_{\mathsf{M}^{-1}}\,{\,\rm d}t+\frac{1}{2}\int_{0}^{T}\big(\|\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E}\|_{\mathsf{M}^{-1}}^{2}+\|{\,\rm d}\mathsf{S}\|_{\mathsf{M}}^{2}\big){\,\rm d}t$
$=-\int_{0}^{T}(\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E},\mathsf{M}{\,\rm d}\mathsf{S})_{\mathsf{M}^{-1}}\,{\,\rm d}t+\frac{1}{2}\int_{0}^{T}(\|\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E}\|_{\mathsf{M}^{-1}}^{2}+\|{\,\rm d}\mathsf{S}\|_{\mathsf{M}}^{2})\,{\,\rm d}t$
$=\frac{1}{2}\int_{0}^{T}\|\partial_{t}\mathsf{z}-\mathsf{L}{\,\rm d}\mathsf{E}-\mathsf{M}{\,\rm d}\mathsf{S}\|_{\mathsf{M}^{-1}}^{2}\,{\,\rm d}t\geq 0,$
where the third equality follows from the first degeneracy condition in (1.6). Thus, $J(\mathsf{z})$ is non-negative for all curves $\mathsf{z}$, and $\mathsf{z}$ is a solution to (1.5) if and only if $J(\mathsf{z})=0$, or equivalently, $J(\mathsf{z})\leq 0$. As shown in [DPZ13], for the Vlasov–Fokker–Planck equation, $J$ is a (multiple) of the large-deviation rate functional from the associated particle systems.
When $\mathsf{E}=0$, the system (1.5) reduces to a gradient flow and the variational characterisation (1.7) corresponds to De Giorgi’s characterisation of gradient flows as curves of maximal slope, see for example [AGS08]. In other words, (1.7) can be viewed as an extension of De Giorgi’s formulation to GENERIC systems. A natural question is whether this can be made rigorous. This was achieved for some simple systems: The kinetic Fokker-Planck equation [DPZ13], a damped harmonic oscillator [JST19, JST21] (these articles use a discretised version of the characterisation (1.7)), and more recently, a fuzzy Boltzmann equation [EH25]. In the following subsection, we discuss previous work on the Landau equation.
1.2 Related work
In Section 3 of this article, we provide a formal calculation showing that the fuzzy Landau equation (1.1) can be cast as a GENERIC system by directly constructing the building blocks $\{\mathsf{M},\,\mathsf{L},\,\mathsf{E},\,\mathsf{S}\}$ and verifying the required conditions in a Schwartz setting, see Appendix A. This formal calculation also applies to the classical Landau equation (1.4), where we take $x=x_{*}$. This means that, formally, the classical Landau equation also has the GENERIC structure. To the best of our knowledge, this has not been demonstrated in the literature.
However, rigorously proving the GENERIC structure of the classical Landau equation (1.4) is a challenging problem. This is because the quadratic form in position $x$, the product $f(x,v)f(x,v_{*})$, in the classical Landau collision operator $Q_{\sf class}(f,f)$ causes the classical collision terms to not naturally be in $L^{1}(\Omega\times\mathbb{R}^{d})$, which is the natural space for the density function. In regards to solvability, within the $L^{1}$-framework, a so called renormalised solution has been studied in [Vil96], among others.
We have introduced a delocalised collisions into the fuzzy Landau equation (1.1), which avoids the above difficulty since
$\|f(x,v)f(x_{*},v_{*})\|_{L^{1}((\Omega\times\mathbb{R}^{d})^{2})}\leq\|f(x,v)\|_{L^{1}(\Omega\times\mathbb{R}^{d})}^{2}.$ (1.8)
This will enable us to rigorously show the GENERIC structure of the fuzzy Landau equation, which sheds new light on these topics.
One closely related model is the spatially homogeneous Landau equation
$\partial_{t}f(v)=Q_{\sf homo}(f,f)(v),\quad f_{t}(v):\mathbb{R}^{d}\to\mathbb{R}_{+}$ (1.9)
with the homogeneous collision term $Q_{\sf homo}$ given in (2.8) below. Recently, Carrillo, Delgadino, Desvillettes and Wu [CDDW24] showed that (1.9) can be interpreted as a gradient flow of the Boltzmann entropy $\mathcal{H}(f)=\int_{\mathbb{R}^{d}}f\log f{\,\rm d}v$ with respect to a certain Wasserstein-type metric.
The weak $L^{1}$ solutions of the homogeneous Landau equation (1.9) have been studied intensively for hard, Maxwellian and soft potentials [Vil98, DV00, DV00a, Wu14, ALL15]. As a consequence of the a priori estimate (1.8), the fuzzy Landau equation (1.1) is structurally similar to the homogeneous Landau equation (1.9). In Part II of this series of articles, we will build on existing results for the homogeneous Landau equation to study the weak $L^{1}$ solutions for the fuzzy Landau equation (1.1).
The Landau equation can also be derived from the grazing limit of the Boltzmann equation, where collisions with small angular deviations play a significant role [Vil98a, He14]. The gradient flow structure of spatially homogeneous Boltzmann equation was studied in [Erb23] and a GENERIC structure for delocalised fuzzy Boltzmann equation was shown by Erbar and the second author in [EH25]. Concerning the grazing limit of the Boltzmann equation, for the spatially homogeneous case, a gradient flow perspective was discussed in [CDW22]. Motivated by these results, in part III [DGH25] we study the grazing limit from the fuzzy Boltzmann equation to the fuzzy Landau equation using their respective GENERIC variational characterisations in [EH25] and the results of this present work.
Delocalised collision models have previously been studied in other contexts, such as the Enskog equation for high-density large diameter particle systems [Ark86, AC89], more general delocalised kinetic models recently discussed in [CCG24], and an optimal transport perspective was discussed in [BMQ25]. Finally, related discrete variational formulations, akin to the celebrated Jordan–Kinderlehrer–Otto [JKO98] scheme, for non-gradient flow systems were studied in [GM12, DPZ14, DT18, ADR22, DAR22], and extensions of symplectic integrators for Hamiltonian systems to GENERIC systems were discussed in [Ött18, SÖ20].
1.3 Variational characterisation
We formally have the following conservation laws for fuzzy Landau equation (1.1)
mass $\int_{\Omega\times\mathbb{R}^{d}}f_{t}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}{\,\rm d}x{\,\rm d}v,$
momentum $\int_{\Omega\times\mathbb{R}^{d}}f_{t}v{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}v{\,\rm d}x{\,\rm d}v,$
energy $\int_{\Omega\times\mathbb{R}^{d}}f_{t}|v|^{2}{\,\rm d}x{\,\rm d}v=\int_{\Omega\times\mathbb{R}^{d}}f_{0}|v|^{2}{\,\rm d}x{\,\rm d}v.$
We define the Boltzmann entropy functional as $\mathcal{H}(f)=\int_{\Omega\times\mathbb{R}^{d}}f\log f{\,\rm d}x{\,\rm d}v$. Formally, the entropy identity holds
$\mathcal{H}(f_{T})-\mathcal{H}(f_{0})=-\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t\leq 0,$
where the entropy dissipation is given by
$\mathcal{D}(f)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{d})^{2}}\kappa ff_{*}\big|\widetilde{\nabla}\log f\big|^{2}{\,\rm d}\eta\geq 0,$ (1.10)
where throughout the paper we use the shortened notation ${\,\rm d}\eta={\,\rm d}x_{*}{\,\rm d}v_{*}{\,\rm d}x{\,\rm d}v$ to denote the Lebesgue measure on $\mathbb{R}^{4d}$ and $\widetilde{\nabla}$ is the fuzzy Landau gradient operator defined below.
A rigorous justification of the conservation laws and entropy identity will be provided in Part II.
As previously mentioned, the fuzzy Landau equation (1.1) can be formulated as a GENERIC system using the building blocks $\{\mathsf{M},\,\mathsf{L},\,\mathsf{E},\,\mathsf{S}\}$ with the verification provided in Appendix A. The energy and entropy are given by
$\mathsf{E}(f)=\int_{\Omega\times\mathbb{R}^{d}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f),$
and the operators $\mathsf{L}$ and $\mathsf{M}$ at $f\in\mathsf{Z}$ are given by
$\mathsf{M}(f)g=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)\quad\text{and}$
$\mathsf{L}(f)g=-\nabla\cdot(f\mathsf{J}\nabla g),\quad\mathsf{J}=\begin{pmatrix}0&\mathsf{id}_{3}\\ -\mathsf{id}_{3}&0\end{pmatrix}$,
where we follow [CDDW24], to define the fuzzy Landau gradient
$\widetilde{\nabla}f(x,v)=\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f(x,v)-\nabla_{v_{*}}f(x_{*},v_{*})\big).$ (1.11)
Note that $A=A(v-v_{*})$ and for the homogeneous Landau equation, $A=|v-v_{*}|^{2+\gamma}$. The corresponding divergence, $\widetilde{\nabla}\cdot$, is defined via integration by parts for a function $B(x,x_{*},v,v_{*}):(\Omega\times\mathbb{R}^{d})^{2}\to\mathbb{R}^{d}$
$\int_{(\Omega\times\mathbb{R}^{d})^{2}}\widetilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\Omega\times\mathbb{R}^{d}}g\widetilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x\quad\forall\,g:\Omega\times\mathbb{R}^{d}\to\mathbb{R}.$ (1.12)
From this, we obtain
$\widetilde{\nabla}\cdot B=\nabla_{v}\cdot\Big(\int_{\Omega\times\mathbb{R}^{d}}\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big(B(x,x_{*},v,v_{*})-B(x_{*},x,v_{*},v)\big){\,\rm d}x_{*}{\,\rm d}v_{*}\Big).$
In this paper, we will rigorously establish the variational characterization (1.7) for the GENERIC structure of the fuzzy Landau equation (1.1) (using $\mathcal{H}$ instead of $\sf S$). We note that the Boltzmann entropy $\mathcal{H}$ defined here, as often in the literature on gradient flows, is negative of the (physical) entropy $\mathsf{S}$ in the GENERIC framework; thus it will be non-increasing along solutions of a GENERIC evolution.
To state the main result, we introduce the transport grazing rate equation
$\partial_{t}f+v\cdot\nabla_{x}f+\frac{1}{2}\widetilde{\nabla}\cdot U_{t}=0,$ (1.13)
where $U:[0,T]\times(\Omega\times\mathbb{R}^{d})^{2}\to\mathbb{R}$ denotes the grazing rate (following the same terminology as in the homogeneous case [CDDW24]). Notice that, when $U_{t}=-\kappa ff_{*}\widetilde{\nabla}\log f$ equation (1.13) recovers the fuzzy Landau equation (1.1). We also define the following curve action associated to (1.13) (see in Definition 3.2 and Lemma 3.3 in Section 3.2.2 for the precise definition and characterisation):
$\mathcal{A}(f,U)=\frac{1}{2}\int_{(\Omega\times\mathbb{R}^{d})^{2}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
The following theorem is the main result of this paper.
Theorem 1.1.
Let the kernels $(A,\kappa)$ satisfy Assumption 2.1. Let $(f_{t},\mathcal{U}_{t})$ be solutions to the transport grazing rate equation (1.13). Let $(f_{t})_{t\in[0,T]}$ be a curve of probability density on $\Omega\times\mathbb{R}^{d}$ such that $\int_{\Omega\times\mathbb{R}^{d}}(\langle x\rangle^{2}+\langle v\rangle^{2+|\gamma|})f_{t}{\,\rm d}x{\,\rm d}v<+\infty$ for any $t\in[0,T]$, initial entropy $|\mathcal{H}(f_{0})|<+\infty$, and satisfies Assumption 2.7 below.
Then we have
$\mathcal{J}_{T}(f,\mathcal{U}):=\mathcal{H}(f_{T})-\mathcal{H}(f_{0})+\frac{1}{2}\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f_{t},\mathcal{U}_{t}){\,\rm d}t\geq 0.$
Moreover, we have $\mathcal{J}_{T}(f,\mathcal{U})=0$ if and only if $f$ is a $\mathcal{H}$-solution for the fuzzy Landau equation (1.1).
The definition of $\mathcal{H}$-solutions is provided in Definition 2.5, following a parallel approach to the $\mathcal{H}$-solutions framework for homogeneous Landau equation introduced by [Vil98a]. The proof of Theorem 1.1, which will be given in Section 3 (it is stated as Theorem 3.16 there), amounts to rigorously verifying a chain rule for the transport grazing rate equation (1.13). We achieve this using regularisation techniques and error estimates. Our analysis requires significant technical improvements compared to the homogeneous case in [CDDW24] to account for the spatial variable $x$ and the transport term $v\cdot\nabla_{x}f(x,v)$. We note that while [CDDW24] is restricted to three dimensions, the variational characterisation in Theorem 1.1 remains valid in all dimensions ($d\geq 2$). Our results also open up a couple of interesting avenues for future research, which will be discussed further in Section 4.
The rest of the paper is organised as follows. In Section 2, we introduce the necessary notations and discuss curves that satisfy Assumption 2.7. In Section 3, we perform regularisations and error estimates to rigorously verify the chain rue and prove our main result, Theorem 1.1. In Section 4, we discuss various GENERIC building blocks for fuzzy Landau equations and possible directions for future works. Finally, some detailed and technical computations are given in Appendices.
This article is the first in a series of works on the fuzzy Landau equation. The second paper [DH25] is devoted to show the existence and regularity results of fuzzy Landau equations. The third paper [DGH25] studies the grazing collision limit of fuzzy Boltzmann equations to fuzzy Landau equations by using the variational method introduced in this paper.
Acknowledgements
M. H. D is funded by an EPSRC Standard Grant EP/Y008561/1. Z. H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.
The authors thank Matthias Erbar for useful comments on this manuscript.
2 Preliminary
We will present the kernel assumptions in Section 2.1. We will introduce the notations in Section 2.2. Assumption 2.7 on the curve in Theorem 1.1 will be discussed in Section 2.3, and related estimates will be provided in Section 2.4.
Throughout this article, for simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
2.1 Kernel assumptions
The spatial kernel $\kappa$ and the interaction kernel
$a(v-v_{*})=A(v-v_{*})\Pi_{(v-v_{*})^{\perp}}$
satisfy the following assumptions.
Assumption 2.1 (Kernel assumptions).
The kernel $A:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is given by
$A(v-v_{*})=|v-v_{*}|^{2+\gamma}\quad\text{with}\quad\gamma\in(-\gamma_{d},1].$ (2.1)
The spatial kernel $k:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is given by
$\kappa(x-x_{*})={k_{1}}\exp(-{k_{2}}\langle x-x_{*}\rangle),$ (2.2)
for some constants ${k_{1},\,k_{2}}>0$ such that $\|\kappa\|_{L^{1}(\mathbb{R}^{d})}=1$.
Remark 2.2.
In Assumption 2.1, we can also consider the spatial kernel $k:\mathbb{R}^{d}\to\mathbb{R}_{+}$ with positive lower and upper bounds
$C_{\kappa}^{-1}\leq\kappa(x-x_{*})\leq C_{\kappa}$ (2.3)
for some constant $C_{\kappa}>0$. Moreover, for a sequence of mollifiers $M_{\beta}$ defined as in (2.9), $\kappa$ satisfies
$C^{-1}C_{\kappa}\leq\big(\kappa^{-1}+\kappa\big)*M_{\beta}\leq CC_{\kappa}\quad\forall\beta\in(0,1)$ (2.4)
for some constant $C>0$. Notice that $\kappa=1$ is included in this case.
Remark 2.3.
In Assumption 2.1, in the hard potential cases, one can also take the kernel $A$ of the following form
$A(v-v_{*})=\overline{A}(v-v_{*})|v-v_{*}|^{2},\quad\forall\,v,\,v_{*}\in\mathbb{R}^{d},$ (2.5)
where $\overline{A}(v-v_{*})=\overline{A}(v_{*}-v)$ and the following bounds hold
$C^{-1}\langle v-v_{*}\rangle^{\gamma+2}\leq\overline{A}(v-v_{*})\leq C\langle v-v_{*}\rangle^{\gamma+2},\quad\gamma\in(-\infty,1]$ (2.6)
for some constant $C>0$.
The spatial kernel $\kappa$ is defined as in (2.2). The pair of kernels $({\overline{A}},\kappa)$ were considered for the fuzzy Boltzmann equation in [EH25]. The fuzzy Landau equation (1.1) with kernel (2.5) is the grazing limit of the fuzzy Boltzmann equation in [EH25] with non-cutoff angular kernels, more details can be found in [DGH25].
Remark 2.4.
The fuzzy Landau equation (1.1) can be seen as an intermediate stage between homogenous (1.9) and inhomogenous (1.4) Landau equations.
We take the spatial kernel $\kappa=\kappa^{\sigma}$, where $\kappa^{\sigma}$ is a renormalised kernel proportional to $\exp\big(-\langle\frac{x}{\sqrt{\sigma}}\rangle\big)$ and $\sigma\in(0,+\infty)$. The kernel $\kappa$ defined in (2.2) can be seen as $\kappa^{\sigma}$ with a fixed $\sigma$.
When $\sigma\to+\infty$, we have $\kappa\to 1$, and we integrate the fuzzy Landau equation (1.1) over $x\in\mathbb{R}^{d}$ at least formally implying the homogeneous Landau equations (1.9). Indeed, we define $F(v)=\int_{\mathbb{R}^{d}}f{\,\rm d}x$, then $F(v)$ satisfies the following homogeneous Landau equations
$\begin{cases}\partial_{t}F=Q_{\sf homo}(F,F),\\ F|_{t=0}=\int_{\mathbb{R}^{d}}f_{0}(x,v){\,\rm d}x,\end{cases}$ (2.7)
where
$Q_{\sf hom}(F,F)=\nabla_{v}\cdot\Big(\int_{\mathbb{R}^{d}}A\big(F(v_{*})\nabla_{v}F(v)-F(v)\nabla_{v_{*}}F(v_{*})\big){\,\rm d}v_{*}\Big).$ (2.8)
When $\sigma\to 0$, the spatial kernel $\kappa^{\sigma}$ converges to a Dirac measure. Formally, the fuzzy Landau equation (1.1) converges to the classical inhomogeneous Landau equation (1.4).
Let $M(z)=C\exp\big(-\langle z\rangle\big)$, where we choose the constant $C>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{d})}=1$. We define the renormalised sequence of mollifiers with parameter $\beta\in(0,1)$
$M_{\beta}(z)=C\beta^{-d}\exp(-\langle{z}/{\beta}\rangle).$ (2.9)
The spatial kernel defined in (2.2) preserves pointwise bounds under the convolution with the sequence (2.9) for all $\beta\in(0,1)$
$\kappa*_{x}M_{\beta}\leq C\kappa$ $\text{and}\quad\kappa^{-1}*_{x}M_{\beta}\leq C\kappa^{-1},$ (2.10)
where $C>0$ is a constant independent of $\beta$.
The proof of (2.10) can be found in [EH25, Lemma 2.1], which is a direct consequence of Peetre’s inequality
$\frac{\langle x\rangle^{p}}{\langle y\rangle^{p}}\leq 2^{|p|/2}\langle x-y\rangle^{|p|}\quad\forall\,x,\,y\in\mathbb{R}^{d},\,\,p\in\mathbb{R},$
which implies
$\big(\langle\cdot\rangle^{p}*M_{\beta}\big)(z)\leq C\langle z\rangle^{p},$ (2.11)
where $C>0$ depends only on $|p|$ and $\|\langle\cdot\rangle^{|p|}M(\cdot)\|_{L^{1}(\mathbb{R}^{d})}$.
2.2 Notations
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
Let $a,\,b\in\mathbb{R}$. We define
$L^{1}_{a,b}(\mathbb{R}^{2d})$ $=\{f\in L^{1}(\mathbb{R}^{2d})\mid\|f\|_{L^{1}_{a,b}(\mathbb{R}^{2d})}<+\infty\},$
with the norms
$\|f\|_{L^{1}_{a,b}(\mathbb{R}^{2d})}\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{a}+\langle v\rangle^{b}\big)f(x,v){\,\rm d}x{\,\rm d}v,$
where we recall that $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese jacket. Let $p,\,q\in[1,\infty)$. The functional space $L^{p}_{v}L^{q}_{x}(\mathbb{R}^{2d})$ consists of all the measure functions $f$ such that
$\|f\|_{L^{p}_{v}L^{q}_{x}}\mathop{=}\limits^{\textrm{def}}\Big(\int_{\mathbb{R}^{d}}\Big(\int_{\mathbb{R}^{d}}|f(x,v)|^{q}{\,\rm d}x\Big)^{\frac{p}{q}}{\,\rm d}v\Big)^{\frac{1}{p}}<+\infty.$
Similarly, we can also define the functional spaces $L^{q}_{x}L^{p}_{v}(\mathbb{R}^{2d})$.
We define $\mathcal{P}(\mathbb{R}^{2d})$ as the space of Borel probability measures on $\mathbb{R}^{2d}$, which is endowed with the weak topology induced by duality with bounded continuous functions in $C_{b}(\mathbb{R}^{2d})$.
For any $a,\,b\in\mathbb{R}$, we define the following subspace of $\mathcal{P}(\mathbb{R}^{2d})$
$\mathcal{P}_{a,b}(\mathbb{R}^{2d})=\{\mu\in\mathcal{P}(\mathbb{R}^{2d})\mid\mathcal{E}_{a,b}(\mu)<+\infty\},$
where
$\mathcal{E}_{a,b}(\mu)\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{2d}}(\langle x\rangle^{a}+\langle v\rangle^{b}){\,\rm d}\mu.$
For functions $f:\mathbb{R}^{2d}\to\mathbb{R}$, we also write
$\mathcal{E}_{a,b}(f)=\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{a}+\langle v\rangle^{b}\big)f(x,v){\,\rm d}x{\,\rm d}v.$
Next we recall the Boltzmann entropy functional for measure $\mu\in\mathcal{P}(\mathbb{R}^{2d})$. If $\mu=f\mathcal{L}$ is absolutely continuous with respect to Lebesgue measure $\mathcal{L}$ with density $f$, we define
$\mathcal{H}(\mu)=\int_{\mathbb{R}^{2d}}f\log f{\,\rm d}x{\,\rm d}v.$
Otherwise, we define $\mathcal{H}(\mu)=+\infty$. In the case of $\mu=f\mathcal{L}$, we also write $\mathcal{H}(f)$. An important property will be used in the subsequent analysis is that the entropy $\mathcal{H}(\mu)$ is weakly lower semi-continuous in the sense that if $\mu_{n}\to\mu$ weakly in $\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{E}_{2,2}(\mu_{n})\to\mathcal{E}_{2,2}(\mu)$ as $n\to\infty$, then
$\mathcal{H}(\mu)\leq\liminf_{n\to\infty}\mathcal{H}(\mu_{n}).$
Throughout this work, $C>0$ denotes a universal constant, which may vary from line to line.
2.3 Assumptions on curves
The fuzzy Landau equation (1.1) can be written as
$(\partial_{t}+v\cdot\nabla_{x})f=\frac{1}{2}\widetilde{\nabla}\cdot\big(\kappa ff_{*}\widetilde{\nabla}\log f\big),$ (2.12)
where we call that the gradient operator $\widetilde{\nabla}$ is defined in (1.11) and the divergence operator $\widetilde{\nabla}\cdot$ is defined via the integration by parts formula (1.12)
$\int_{\mathbb{R}^{4d}}\widetilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\mathbb{R}^{2d}}g\widetilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x$
for all $B(x,x_{*},v,v_{*}):\mathbb{R}^{4d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}^{2d}\to\mathbb{R}$. We follow [Vil98a] to define a $\mathcal{H}$-solution for fuzzy Landau equation (1.1).
Definition 2.5 ($\mathcal{H}$-solution).
Let $\gamma\in(-\gamma_{d},1]$ and $A(v-v_{*})=|v-v_{*}|^{2+\gamma}$ is defined as in (2.1). Let
$f\in C([0,T];\mathscr{D}^{\prime}(\mathbb{R}^{2d}))\cap L^{1}([0,T];L^{1}_{2,2+\gamma_{+}}(\mathbb{R}^{2d})),$
where $\gamma_{+}=\max\{\gamma,0\}$. We say $f$ is a $\mathcal{H}$-solution of (1.1) if
$f_{t}\geq 0\quad\text{and}\quad\int_{\mathbb{R}^{2d}}f_{t}(x,v){\,\rm d}x{\,\rm d}v=1\quad\forall\,t\in[0,T];$
The entropy dissipation is time integrable
$\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\log f|^{2}{\,\rm d}\eta{\,\rm d}t<+\infty;$
And the following identity holds
$\int_{\mathbb{R}^{2d}}f_{0}\varphi(0){\,\rm d}x{\,\rm d}v+\int_{0}^{T}\int_{\mathbb{R}^{2d}}f(\partial_{t}\varphi+v\cdot\nabla_{x}\varphi){\,\rm d}x{\,\rm d}v{\,\rm d}t$ (2.13)
$=-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}\varphi\cdot\widetilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t$
for all $\varphi\in C^{\infty}_{c}([0,T)\times\mathbb{R}^{2d})$.
Notice that, on the domain $\mathbb{T}^{d}\times\mathbb{R}^{d}$, we do not need an extra moment assumption on $x$.
Remark 2.6.
  • The right-hand side of (2.13) is well-defined. Indeed, by the Cauchy–Schwartz inequality,
    $\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}\varphi\cdot\widetilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t\Big|$
    $\leq$ ${}\Big|\int_{0}^{T}2\mathcal{D}(f){\,\rm d}t\Big|^{\frac{1}{2}}\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big|^{\frac{1}{2}}.$
    In the case of $\gamma\in[-2,1]$, we have
    $\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}\leq C(\varphi,\gamma)ff_{*}(|v|^{2+\gamma}+|v_{*}|^{2+\gamma})\in L^{1}([0,T]\times\mathbb{R}^{4d}).$
    In the case of very soft potential $\gamma\in(-\gamma_{d},-2)$, we assume that $\operatorname{supp}(\varphi)\subset\mathbb{R}^{d}_{x}\times B^{R}_{v}$. By using Hölder and Young’s inequalities, we have
    $\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big|$ (2.14)
    $\leq$ ${}C(\varphi)\int_{0}^{T}\int_{\mathbb{R}^{2d}\times(B^{R})^{2}}ff_{*}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t$
    $\lesssim$ ${}\|f\|_{L^{\infty}([0,T];L^{1}(\mathbb{R}^{2d})}\|f\|_{L^{1}([0,T];L^{1}_{x}L^{p}(B^{R}_{v})}\||v|^{\gamma+2}\|_{L^{p^{\prime}}(B^{2R})}.$
    where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. The Fisher information estimate in Lemma 2.9 and the Sobolev embedding ensure that $f\in L^{1}([0,T];L^{1}_{x}(L^{k}_{v})_{\operatorname{loc}}$, where $k=\frac{d}{d-2}$ for $d\geq 3$, and $k\in(1,+\infty)$ for $d=2$. We choose $(p,p^{\prime})=(k,d/2)$ in (2.14), then $|\gamma+2|p^{\prime}<d$ ensures that $|v|^{\gamma+2}\in L^{p^{\prime}}_{\operatorname{loc}}(\mathbb{R}^{d})$. The detailed discussion can be found in Part II.
  • In the homogeneous Landau cases in [Vil98a], one can justify the $\mathcal{H}$-solutions in a different way. In the homogeneous case with $\psi=\psi(v)\in C^{\infty}_{c}(\mathbb{R}^{d})$, we have
    $\widetilde{\nabla}_{\sf homo}\psi(v)=|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})\big),$
    which satisfies
    $|\widetilde{\nabla}_{\sf homo}\psi|$ $\leq|v-v_{*}|^{1+\frac{\gamma}{2}}|\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})|$
    $\leq|v-v_{*}|^{2+\frac{\gamma}{2}}\|\psi\|_{\operatorname{Lip}(\mathbb{R}^{d})},$
    where $2+\frac{\gamma}{2}>0$. However, this is not the case for fuzzy Landau equations, since we also need to take into account the distance between $x$ and $x_{*}$ in
    $\nabla_{v}\varphi(x,v)-\nabla_{v_{*}}\varphi(x_{*},v_{*}).$
  • The entropy dissipation bound is natural since we will show in Part II that the following entropy inequalities hold
    $\mathcal{H}(f_{T})-\mathcal{H}(f_{0})+\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t\leq 0.$
    We note that $f\in L^{1}_{2,2}(\mathbb{R}^{2d})$ ensures $\mathcal{H}(f_{T})>$-$\infty$.
In the main Theorem 1.1, we consider the curves $\mu_{t}=f_{t}\mathcal{L}$ with the following assumptions according to the three cases of kernels in Assumption 2.1.
Assumption 2.7.
For $(f_{t}\mathcal{L})_{t\in[0,T]}\subset\mathcal{P}_{0,2+|\gamma|}(\mathbb{R}^{2d})$, we assume
$f\in L^{1}\big([0,T];L^{\infty}_{2,0}(\mathbb{R}^{2d})\big)\cap L^{\infty}\big([0,T];L^{\infty}_{0,2+|\gamma|}(\mathbb{R}^{2d})\big).$
In addition, in the soft potential case $\gamma\in(-\gamma_{d},0)$, we assume
$\langle v\rangle^{2+\gamma}f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x});$
In the hard potential case $\gamma\in(0,1]$, we assume
$f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$
for some $p\in\big(\frac{d}{d-|\gamma|},+\infty\big)$.
Remark 2.8.
One can make a similar assumption by replacing $L^{p}_{v}L^{1}_{x}$ by $L^{1}_{x}L^{p}_{v}$. The $L^{p}$ bound is to ensure the integrability of $\int_{\mathbb{R}^{4d}}ff_{*}|v-v_{*}|^{\gamma}{\,\rm d}\eta$, for which we use in Lemma 2.10 below, and we make no distinction between $L^{1}_{x}L^{p}_{v}$ and $L^{p}_{v}L^{1}_{x}$ there.
In the follow-up Part II paper, we will show the existence and regularity of the $\mathcal{H}$ solutions to the fuzzy Landau equation (1.1).
Concerning the propagation of moment and $L^{p}_{v}L^{1}_{x}$ norms, we first see the case of the spatial kernel $\kappa=1$. We integrate (1.1) over $x\in\mathbb{R}^{d}$ and define $F(v)=\int_{\mathbb{R}^{d}}f{\,\rm d}x$, which satisfies the homogeneous Landau equations (2.7). The regularity results for homogeneous Landau equations hold: For $\gamma\in[-2,1]$, all moment and $L^{p}$ norms for $p>3$ propagated, see [DV00] and [Wu14] for hard and soft potential cases; For the very soft potential case, the moment and $L^{p}$ norms propagated only known to be local in time or global in time for small initial data, see for example [ALL15].
Concerning the variable spatial kernel $\kappa$, similar results hold on torus $\mathbb{T}^{d}=\mathbb{R}^{d}/{\mathbb{Z}}^{d}$ with $L^{p}_{v}L^{1}_{x}$ replaced by $L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})$. The proof will be contained in the upcoming Part II. Notice that $L^{p}(\mathbb{R}^{d};L^{1}(\mathbb{T}^{d}))\subset L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})$ since by Minkowski and Hölder inequalities
$\Big(\int_{\mathbb{R}^{d}}\Big(\int_{\mathbb{T}^{d}}f{\,\rm d}x\Big)^{p}{\,\rm d}v\Big)^{\frac{1}{p}}\leq\int_{\mathbb{T}^{d}}\Big(\int_{\mathbb{R}^{d}}f^{p}{\,\rm d}v\Big)^{\frac{1}{p}}{\,\rm d}x\leq\|f\|_{L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})}.$
2.4 Entropy dissipation and Fisher information
In this Subsection, we consider the curves satisfying Assumption 2.7, and the kernel $(A,\kappa)$ satisfying Assumption 2.1.
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
In the subsequent analysis, we will use the following identity
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=\frac{1}{2}\sum_{i,j=1}^{d}|x_{i}y_{j}-x_{j}y_{i}|^{2}\quad\forall\,x,\,y\in\mathbb{R}^{d},$
which can be verified by direct computations:
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)$ $=|x|^{2}\Big[y-\frac{x\otimes x}{|x|^{2}}y\Big]\cdot y$
$=|x|^{2}|y|^{2}-(x\otimes x)y\cdot y$
$=|x|^{2}|y|^{2}-\sum_{i,j}(x\otimes x)_{ij}y_{i}y_{j}$
$=(\sum_{i}x_{i}^{2})(\sum_{j}y_{j}^{2})-\sum_{i,j}x_{i}x_{j}y_{i}y_{j}$
$=(\sum_{i}x_{i}^{2})(\sum_{j}y_{j}^{2})-\Big(\sum_{i}x_{i}y_{i}\Big)^{2}$
$=\frac{1}{2}\sum_{ij}|x_{i}y_{j}-x_{j}y_{i}|^{2}.$
Notice that in dimension three, this reads $|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=|x\times y|^{2}$. To simplify the notation, we write
$[x,y]_{ij}=x_{i}y_{j}-x_{j}y_{i},\quad i,j=1,\dots,d.$
Then the entropy dissipation (1.10) can be written as
$\mathcal{D}(f)=$ $\frac{1}{4}\sum_{i,j=1}^{d}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|v-v_{*}|^{\gamma}|[v-v_{*},\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}]_{ij}\big|^{2}{\,\rm d}\eta.$ (2.15)
Indeed, we have
$|\widetilde{\nabla}\log f|^{2}$ (2.16)
$=$ ${}|v-v_{*}|^{\gamma}|v-v_{*}|^{2}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)\cdot\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)$
$=$ ${}\frac{|v-v_{*}|^{\gamma}}{2}\sum_{i,j=1}^{d}\big|[v-v_{*},\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}]_{ij}\big|^{2}.$
We define the Fisher and cross-Fisher information, respectively, by
$\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\frac{\nabla_{v}f}{f}\Big|^{2}{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\sum_{i,j=1}^{d}\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\Big[v,\frac{\nabla_{v}f}{f}\Big]_{ij}\Big|^{2}{\,\rm d}x{\,\rm d}v.$
We will follow [Des20] and [Des16] of homogeneous Landau equations to show that the Fisher and cross-Fisher information can be bounded by the entropy dissipation.
Lemma 2.9.
Let $f\in\mathcal{P}\cap L\log L\cap\mathcal{S}(\mathbb{R}^{2d})$. Let $\mathcal{D}(f)$ be the entropy dissipation defined as in (2.15), where $\gamma\in(-\gamma_{d},1]$ and the kernels satisfy
$A(v-v_{*})=|v-v_{*}|^{2+\gamma}.$
The following bounds hold
$\sum_{i,j=1}^{d}\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\Big[\frac{\nabla_{v}f}{f},v\Big]_{ij}\Big|^{2}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\frac{\nabla_{v}f}{f}\Big|^{2}{\,\rm d}x{\,\rm d}v$ $\leq C(f)\big(1+\mathcal{D}(f)\big).$ (2.17)
  • Let $\gamma\in(-\gamma_{d},0]$. Then we have
    $C(f)\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{7}+\|f\|_{L^{1}_{0,|\gamma|}(\mathbb{R}^{2d})}.$
  • Let $\gamma\in(0,1]$. Let $p>\frac{d}{d-\gamma}$. Then we have
    $C(f)\lesssim$ $\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\big(\|f\|_{L^{1}_{0,2+\gamma}(\mathbb{R}^{2d})}+\|f\|_{L^{1}_{0,\gamma}(\mathbb{R}^{2d})}$
    $\quad+\min(\|\langle v\rangle^{\gamma}f\|_{L^{p}_{v}L^{1}_{x}},\|\langle v\rangle^{\gamma}f\|_{L^{1}_{x}L^{p}_{v}})\big).$
Proof.
The proof follows [Des16, Corollary 2.7] with appropriate modification according to the spatial kernel $x$.
We define $\partial_{j}\mathop{=}\limits^{\textrm{def}}\partial_{v_{j}}$ and for $i,j=1,\ldots,d$:
$q_{ij}=q_{ij}^{f}(x,v,y,w)=\Big[v-w,\frac{\nabla f}{f}(x,v)-\frac{\nabla f}{f}(y,w)\Big]_{i,j}$ (2.18)
$=(v_{i}-w_{i})\Big(\frac{\partial_{j}f}{f}(x,v)-\frac{\partial_{j}f}{f}(y,w)\Big)-(v_{j}-w_{j})\Big(\frac{\partial_{i}f}{f}(x,v)-\frac{\partial_{i}f}{f}(y,w)\Big)$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)+\Big(w_{j}\frac{\partial_{i}f}{f}(x,v)-w_{i}\frac{\partial_{j}f}{f}(x,v)\Big)+\Big(v_{j}\frac{\partial_{i}f}{f}(y,w)-v_{i}\frac{\partial_{j}f}{f}(y,w)\Big)$
$\quad+\Big(w_{i}\frac{\partial_{j}f}{f}(x,v)-w_{j}\frac{\partial_{i}f}{f}(x,v)\Big).$
We write $q^{f}=(q_{ij})_{i,j}$. The entropy dissipation (2.15) can be written as
$\mathcal{D}(f)=\frac{1}{4}\int_{\mathbb{R}^{4d}}\kappa(x-y)|v-w|^{\gamma}f(x,v)f(y,w)|q^{f}|^{2}{\,\rm d}\eta.$ (2.19)
We have the following identity:
$\Big[v,\frac{\nabla_{v}f}{f}(x,v)\Big]_{ij}-w_{i}\frac{\partial_{i}f}{f}(x,v)+w_{j}\frac{\partial_{j}f}{f}(x,v)$ (2.20)
$=$ ${}v_{i}\frac{\partial_{j}f}{f}(y,w)-v_{j}\frac{\partial_{i}f}{f}(y,w)-\Big(w_{i}\frac{\partial_{j}f}{f}(y,w)-w_{j}\frac{\partial_{i}f}{f}(y,w)\Big)+q_{ij}.$
To obtain a better bound, we define the weight function $\phi(r)=(1+2r)^{-\frac{|\gamma|}{4}-\frac{1}{2}}$ and we write $\phi=\phi(|w|^{2}/2)=\langle w\rangle^{-\frac{|\gamma|}{2}-1}$. Notice that $|\phi|\leq 1$, $|\phi^{\prime}(r)|\leq\frac{3}{2}$ and $\nabla\phi=w\phi^{\prime}(|w|^{2}/2)$.
We multiply (2.18) by $\phi\kappa(x-y)f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{2d}$, noting that this will annihilate the last bracket term in the second line of (2.18), to derive
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{i}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w-v_{j}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w.$ (2.21)
Next we multiply (2.18) by $\phi\kappa w_{i}f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{2d}$ to derive
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}^{2}f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{i}\int_{\mathbb{R}^{2d}}\phi^{\prime}w_{i}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$-v_{j}\int_{\mathbb{R}^{2d}}(\phi+w_{i}^{2}\phi^{\prime})\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w,$ (2.22)
where to obtain the last term, we have used the following calculations
$\int_{\mathbb{R}^{2d}}\Big(w_{i}\frac{\partial_{j}f}{f}(y,w)-w_{j}\frac{\partial_{i}f}{f}(y,w)\Big)w_{i}\phi\kappa f(y,w)\,dydw$
$\qquad=\int_{\mathbb{R}^{2d}}w_{i}^{2}\phi\kappa\partial_{j}f(y,w){\,\rm d}y{\,\rm d}w-\int_{\mathbb{R}^{2d}}w_{i}w_{j}\phi\kappa\partial_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=-\int_{\mathbb{R}^{2d}}w_{i}^{2}w_{j}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}w_{j}\partial_{i}(w_{i}\phi)\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=-\int_{\mathbb{R}^{2d}}w_{i}^{2}w_{j}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}w_{j}(\phi+w_{i}^{2}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=\int_{\mathbb{R}^{2d}}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w.$
Similarly, testing (2.18) by $\phi\kappa w_{j}f(y,w)$ yields
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}^{2}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$-v_{j}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{i}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{j}\int_{\mathbb{R}^{2d}}(\phi\kappa+w_{j}^{2}\phi^{\prime}\kappa)f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w.$ (2.23)
From (2.21), (2.22) and (2.23), we obtain a $3\times 3$-system of linear equations for $v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)$, $\frac{\partial_{i}f}{f}(x,v)$ and $\frac{\partial_{j}f}{f}(x,v)$. Solving this system by Cramer’s formula, we get
$\frac{\partial_{i}f}{f}(x,v)=\frac{\Delta_{i}(f)}{\Delta_{\phi}(f)}\quad\text{and}\quad v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)=\frac{\Delta_{ij}(f)}{\Delta_{\phi}(f)}.$
In the above equation: $\Delta_{i}(f)$ and $\Delta_{ij}(f)$ are given by
$\Delta_{i}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}1&q_{ij}+P^{ij}_{1}&w_{i}\\ w_{i}&w_{i}q_{ij}+P^{ij}_{2}&w_{i}^{2}\\ w_{j}&w_{j}q_{ij}+P^{ij}_{3}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big),$
$\Delta_{ij}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}q_{ij}+P^{ij}_{1}&w_{i}&w_{j}\\ w_{i}q_{ij}+P^{ij}_{2}&w_{i}w_{j}&w_{i}^{2}\\ w_{j}q_{ij}+P^{ij}_{3}&w_{j}^{2}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big),$
where
$P^{ij}_{1}$ $=(v_{j}w_{i}-v_{i}w_{j})\frac{\phi^{\prime}}{\phi},$
$P^{ij}_{2}$ $=v_{j}-w_{j}+(v_{j}w_{i}^{2}-v_{i}w_{i}w_{j}))\frac{\phi^{\prime}}{\phi},$
$P^{ij}_{3}$ $=w_{i}-v_{i}+(v_{j}w_{i}w_{j}-v_{i}w_{j}^{2})\frac{\phi^{\prime}}{\phi}.$
And $\Delta_{\phi}(f)$ is given by
$\Delta_{\phi}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}1&w_{i}&w_{j}\\ w_{i}&w_{i}w_{j}&w_{i}^{2}\\ w_{j}&w_{j}^{2}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big).$
We define
$\mathbb{F}(x,w)=\int_{\mathbb{R}^{d}}f(y,w)\kappa(x-y){\,\rm d}y.$
By using the convexity of $x\mapsto x\log x$ and Jensen’s inequality, we have $\mathcal{H}(\mathbb{F})\leq\mathcal{H}(f)$. And we have $\|\mathbb{F}\|_{L^{1}_{2,2}(\mathbb{R}^{2d})}\leq C(\|\kappa\|_{L^{\infty}})\|f\|_{L^{1}_{2,2}(\mathbb{R}^{2d})}$. One can closely follow [Des16, Proposition 4] to show that $\mathbb{F}\in L^{1}_{2,2}(\mathbb{R}^{2d})$ and $|\mathcal{H}(\mathbb{F})|<+\infty$ ensure that $\Delta_{\phi}\geq C>0$. Then we have
$\Big|\frac{\nabla_{v}f}{f}(x,v)\Big|\lesssim\sum_{i,j}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\int_{\mathbb{R}^{2d}}\phi\kappa f(\sum_{k=1}^{3}|P^{ij}_{k}|+|q_{ij}|\langle w\rangle\big){\,\rm d}y{\,\rm d}w$
$\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\Big(\langle v\rangle\int_{\mathbb{R}^{2d}}f\langle w\rangle^{2}(\phi+\phi^{\prime})\kappa{\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa f|q|\langle w\rangle{\,\rm d}y{\,\rm d}w\Big)$
$\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\Big(\|\kappa\|_{L^{\infty}}\langle v\rangle\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}+\int_{\mathbb{R}^{2d}}\kappa f|q|{\,\rm d}y{\,\rm d}w\Big).$
Using (2.20), we deduce that the same bounds also hold for $|\Big[v,\frac{\nabla_{v}f}{f}\Big]_{i,j}|$.
Hence, by using Cauchy–Schwarz inequality, we have
$\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}f(x,v)\Big|\frac{\nabla_{v}f}{f}(x,v)\Big|^{2}{\,\rm d}x{\,\rm d}v$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\Big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}$
$+\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}f(x,v)\big|\int_{\mathbb{R}^{2d}}\kappa f(y,w)|q|{\,\rm d}y{\,\rm d}w\big|^{2}{\,\rm d}x{\,\rm d}v\Big)$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\Big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}$
$+\int_{\mathbb{R}^{2d}}f(x,v)\big(\int_{\mathbb{R}^{2d}}\kappa f(y,w)|q|^{2}|v-w|^{\gamma}{\,\rm d}y{\,\rm d}w\big)$
$\times\Big(\int_{\mathbb{R}^{2d}}\kappa f(y,w)|v-w|^{-\gamma}\langle v\rangle^{\gamma}\Big){\,\rm d}x{\,\rm d}v\Big)$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}+\mathcal{D}(f)C^{\pm}_{-\gamma,0}(f)\big)$
where $C^{\pm}_{-\gamma,0}$ are given in (2.25) and (2.24) below. The same bounds hold for the cross Fisher information.
The following lemma shows the bounds on the convolution $\|f*_{v}|v-\cdot|^{\alpha}\|_{L^{1}_{x}}$, which are used in Lemma 2.9.
Lemma 2.10.
Let $\alpha\in(-d,\infty)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{2d})$. The following estimates hold:
  • If $\alpha\in[0,\infty)$, then
    $\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\leq C^{+}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{d},$
    where
    $C^{+}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{\alpha+\delta_{+}}f\|_{L^{1}(\mathbb{R}^{2d})},\quad\delta_{+}=\max\{\delta,0\}.$ (2.24)
  • If $\alpha\in(-d,0)$, then we have
    $\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\leq C^{-}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{d},$
    where
    $C^{-}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{1}(\mathbb{R}^{2d})}+\min(\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{p}_{v}L^{1}_{x}},\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{1}_{x}L^{p}_{v}})$ (2.25)
    for some $p>\frac{d}{d-|\alpha|}$.
Proof.
The proof follows [CDDW24] with appropriate modification according to the spatial kernel $x$.
  • The case of $\alpha\in[0,+\infty)$: We apply straightforwardly $|v-v_{*}|^{\alpha}\lesssim|v|^{\alpha}+|v_{*}|^{\alpha}$.
  • The case of $\alpha\in(-d,0)$: We consider the cases of $\{|v|\leq 1\}$ and $\{|v|\geq 1\}$ separately.
    In the case of $|v|\leq 1$, we only need to show the boundedness of $\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{d})}$. Notice that
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.26)
    $=$ ${}\Big(\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*},$
    where we have used the inequality $|v-v_{*}|^{\alpha}\leq 1$ (since $|v-v_{*}|\geq 1$ and $\alpha<0$) to estimate the first term. On the domain $\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}$, we have
    $\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.27)
    $=\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}\chi_{B_{1}}|v-v_{*}|^{\alpha}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=\big(\int_{\mathbb{R}^{d}}\langle\cdot\rangle^{\delta}f(x_{*},\cdot){\,\rm d}x_{*}\big)*_{v_{*}}(\chi_{B_{1}}|\cdot|^{\alpha})$
    $\leq\|\langle v\rangle^{\delta}f\|_{L^{p}_{v}L^{1}_{x}}\||v|^{\alpha}\|_{L^{p^{\prime}}(B_{1})},$
    where $B_{1}=\{v\mid|v|\leq 1\}$ and $\chi_{B_{1}}$ denotes its characteristic function, and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Notice that $p>\frac{d}{d-|\alpha|}$ ensures that $0\leq|\alpha|p^{\prime}<d$, and hence $|v|^{\alpha}\in L^{p^{\prime}}(B_{1})$. Notice that, in the above, if we first apply Young’s inequality then integrate of $x_{*}$, then we obtain the bound in terms of $\|\langle v\rangle^{\delta}f\|_{L^{1}_{x}L^{p}_{v}}$. This argument also applies to the following estimates.
    In the case of $|v|\geq 1$, we split the integral region as following
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=$ ${}\Big(\int_{\mathbb{R}^{d}\times\{|v_{*}|\leq\frac{|v|}{2}\}}+\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\Big(\frac{|v|}{2}\Big)^{\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\leq\frac{|v|}{2}\}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $+\Big(\frac{|v|}{2}\Big)^{\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}2^{-\alpha}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+2^{-\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}.$
    Note that to obtain the last inequality above, we have used the fact that $|v|\geq 1$ and $\alpha<0$. Similar to (2.27), on the domain $\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}$, we have
    $\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\Big(\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\big(\int_{\mathbb{R}^{d}}f(x_{*},\cdot)\langle\cdot\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}\big)*_{v}(\chi_{B_{1}}|\cdot|^{\alpha})$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}\||v|^{-\gamma}\|_{L^{p^{\prime}}(B_{1})}.$
    We combine all the cases and conclude that
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\lesssim$ ${}\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}.$
Corollary 2.11.
Let $\alpha\in(-d,0)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{2d})$. Then we have
$\sup_{v\in\mathbb{R}^{d}}\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\lesssim\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+\min(\|\langle v\rangle^{\delta}f\|_{L^{1}_{x}L^{p}_{v}},\|\langle v\rangle^{\delta}f\|_{L^{p}_{v}L^{1}_{x}})$
for some $p>\frac{d}{d-|\alpha|}$.
Proof.
We repeat (2.26)-(2.27) in the proof of Lemma 2.10. ∎
3 Variational characterisation
In this section, we will first introduce and discuss the transport grazing equation in Section 3.1. We will prove a chain rule for transport grazing rate equations in Section 3.2. We prove a variational characterisation Theorem 3.16 (which was stated as Theorem 1.1 in the Introduction) for the fuzzy Landau equation (1.1) in Section 3.3.
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to $\mathbb{T}^{d}\times\mathbb{R}^{d}$ and $\mathbb{R}^{2d}$.
3.1 Transport grazing rate equation
We consider curves $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}(\mathbb{R}^{2d})$ derived by transport and grazing mechanism via a transport grazing rate continuity equation
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0,$ (3.1)
where $\mathcal{U}_{t}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ denotes the grazing rate, and $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ denotes the space of vector-valued signed Borel measures with bounded total variation on $\mathbb{R}^{4d}$. The space $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ is endowed with the weak* topology induced by $C_{0}(\mathbb{R}^{4d};\mathbb{R}^{d})$ with continuous functions vanishing at infinity. We recall again that the gradient operator, $\widetilde{\nabla}$, and the associated divergence operator, $\widetilde{\nabla}\cdot$, are defined, respectively, in (1.11) and (1.12). Also, we have denoted
${\,\rm d}\eta={\,\rm d}x{\,\rm d}v{\,\rm d}x_{*}{\,\rm d}v_{*}$
the Lebesgue measure on $\mathbb{R}^{4d}$.
Definition 3.1.
We say $(\mu_{t},\mathcal{U}_{t})$ is a pair of solutions to the transport grazing rate equation (3.1) if
  1. 1.
    $\mu_{t}:[0,T]\to\mathcal{P}(\mathbb{R}^{2d})$ is weakly continuous;
  2. 2.
    $(\mathcal{U}_{t})_{t\in[0,T]}$ is a family of Borel measures in $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$;
  3. 3.
    $\int_{0}^{T}{\,\rm d}|\mathcal{U}_{t}|(\mathbb{R}^{4d})<\infty$;
  4. 4.
    For any $\varphi\in C^{\infty}_{c}(\mathbb{R}^{2d})$ the following equality holds
    $\frac{d}{dt}\int_{\mathbb{R}^{2d}}\varphi{\,\rm d}\mu_{t}-\int_{\mathbb{R}^{2d}}v\cdot\nabla\varphi{\,\rm d}\mu_{t}=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\varphi{\,\rm d}\mathcal{U}_{t}.$
    We define the set of such pairs $(\mu_{t},\mathcal{U}_{t})$ by $\mathcal{TGRE}_{T}$.
If the pair $(\mu_{t},\mathcal{U}_{t})$ has densities $f_{t}$ and $U_{t}$ with respect to Lebesgue measure on $\mathbb{R}^{2d}$ and $\mathbb{R}^{4d}$, we will also write $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$.
We define the function $\alpha:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}$ by letting
$\alpha(s,u)=\left\{\begin{aligned} &\frac{u^{2}}{2s},\quad s\neq 0,\\ &0,\quad u=0,\,s=0,\\ &+\infty,\quad u\neq 0,\,s=0.\end{aligned}\right$. (3.2)
Notice that the function $(a,b)\mapsto\frac{a^{2}}{b}$ is joint convex on $\mathbb{R}\times\mathbb{R}_{+}$. The function $\alpha$ is lower semicontinuous, convex and 1-homogeneous, i.e. $\alpha(rs,ru)=r\alpha(s,u)$ for all $r\geq 0$.
Next we define the action functional for $\mu\in\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$.
Definition 3.2 (action functional).
We define the action functional
$\mathcal{A}(\mu,\mathcal{U})=\int_{\mathbb{R}^{4d}}\alpha\Big(\frac{{\,\rm d}\mu_{1}}{{\,\rm d}\lambda},\frac{{\,\rm d}\mathcal{U}}{{\,\rm d}\lambda}\Big){\,\rm d}\lambda,$ (3.3)
where
$\mu_{1}({\,\rm d}\eta)\mathop{=}\limits^{\textrm{def}}\mu_{t}({\,\rm d}x{\,\rm d}v)\mu_{t}({\,\rm d}x_{*}{\,\rm d}v_{*})\quad\text{and}\quad\lambda\mathop{=}\limits^{\textrm{def}}\mu_{1}+|\mathcal{U}_{t}|.$
Notice that $\mu_{1}\in\mathcal{P}(\mathbb{R}^{4d})$ and $\mathcal{A}(\mu,\mathcal{U})$ is well-defined since $\alpha$ is 1-homogeneous.
The following lemma shows that if $\mu$ has a density with respect to Lebesgue measure $\mathcal{L}$, then the bounded action functional ensures the existence of density for $\mathcal{U}$.
Lemma 3.3.
If $\mu=f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ and $\mathcal{A}(\mu,\mathcal{U})<+\infty$, then there exists a Borel measure $M:\mathbb{R}^{4d}\to\mathbb{R}^{d}$ such that $\mathcal{U}=Mff_{*}\kappa{\,\rm d}\eta=U{\,\rm d}\eta$ and
$\mathcal{A}(\mu,\mathcal{U})=\frac{1}{2}\int_{\mathbb{R}^{4d}}|M|^{2}ff_{*}\kappa{\,\rm d}\eta=\frac{1}{2}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
Proof.
The proof straightforwardly follows [Erb23] and [CDDW24]. Let ${\,\rm d}\widetilde{\lambda}=\kappa{\,\rm d}\lambda$, ${\,\rm d}\widetilde{\mu}_{1}=\kappa{\,\rm d}\mu_{1}$ and ${\,\rm d}\widetilde{\mathcal{U}}=\kappa{\,\rm d}\mathcal{U}$. Then 1-homogeneity of $\alpha$ ensures that
$\mathcal{A}(\mu_{t},\mathcal{U}_{t})=\int_{\mathbb{R}^{4d}}\alpha\Big(\frac{{\,\rm d}\widetilde{\mu}_{1}}{{\,\rm d}\widetilde{\lambda}},\frac{{\,\rm d}\widetilde{\mathcal{U}}}{{\,\rm d}\widetilde{\lambda}}\Big){\,\rm d}\widetilde{\lambda}$$\infty$. (3.4)
It is sufficient to show that $\widetilde{\mathcal{U}}$ is absolutely continuous with respect to $\widetilde{\mu}_{1}$. Let $S\subset\mathbb{R}^{4d}$ such that $\widetilde{\mu}_{1}(S)=0$, then $\frac{{\,\rm d}\widetilde{\mu}_{1}}{{\,\rm d}\widetilde{\lambda}}=0$ $\widetilde{\lambda}-$almost everywhere. By definition of $\alpha$ in (3.2) and the finiteness (3.4), we have $\frac{{\,\rm d}\widetilde{\mathcal{U}}}{{\,\rm d}\widetilde{\lambda}}=0$ $\widetilde{\lambda}-$almost everywhere. Hence, $\mathcal{U}$ is absolutely continuous with respect to $\widetilde{\mu}_{1}$ and $\sigma$. ∎
For the pair $(\mu,\mathcal{U})$ with densities $(f,U)$ with respect to Lebesgue measures $\mathcal{L}$, we also write $\mathcal{A}(f,U)\mathop{=}\limits^{\textrm{def}}\mathcal{A}(\mu,\mathcal{U})$.
Lemma 3.4 (Lower semicontinuity).
Let $\mu_{n}\rightharpoonup\mu$ in $\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}_{n}\overset{*}{\rightharpoonup}\mathcal{U}$ in $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$. We have
$\mathcal{A}(\mu,\mathcal{U})\leq\liminf_{n\to\infty}\mathcal{A}(\mu_{n},\mathcal{U}_{n})\quad\text{and}\quad\mathcal{D}(\mu,\mathcal{U})\leq\liminf_{n\to\infty}\mathcal{D}(\mu_{n},\mathcal{U}_{n}).$
Proof.
This lemma is a direct consequence of [But89, Theorem 3.4.3]. ∎
We have the following integrability estimate on the grazing rate.
Lemma 3.5 (Integrability estimates).
Let $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$. If
$C_{A}\mathop{=}\limits^{\textrm{def}}\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t<+\infty\quad\text{and}\quad C_{E}=\int_{0}^{T}\mathcal{E}_{2,2+|\gamma|}(f_{t}){\,\rm d}t<+\infty,$
then for $\gamma\in(-\gamma_{d},1]$, we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big){\,\rm d}\eta{\,\rm d}t\leq\sqrt{C_{A}C_{E}}.$ (3.5)
When $\gamma\in(-\gamma_{d},-2)$, if in addition $f\in L^{1}([0,T];L^{p}_{v}L^{1}_{x})$ for some $p>\frac{d}{d-|2+\gamma|}$, then we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t\leq\sqrt{CC_{A}},$ (3.6)
where $C\lesssim\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T]\times\mathbb{R}^{2d})}+\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T];L^{p}_{v}L^{1}_{x})}$.
Proof.
We first show (3.5). By Cauchy–Schwarz inequality, we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle\big){\,\rm d}\eta{\,\rm d}t$ (3.7)
$\leq\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}\times$
$\quad\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle^{2+|\gamma|}+\langle x\rangle^{2}\big){ff_{*}}\kappa{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}$
$=\sqrt{C_{A}C_{E}}.$
Concerning (3.6), similarly to (3.7), we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}r$
$\leq{C_{A}}^{\frac{1}{2}}\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}.$
By using of Corollary 2.11, we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\lesssim 1+\|f\|_{L^{1}([0,T];L^{p}_{v}L^{1}_{x})}.$
3.2 Chain rule
In this section, we prove the following chain rule for the transport grazing rate equation
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0.$
Theorem 3.6 (Chain rule).
Let $\gamma\in(-\gamma_{d},1]$. Let $(\mu_{t},\mathcal{U}_{t})\in\mathcal{TGRE}_{T}$ such that $|\mathcal{H}(\mu_{0})|<+\infty$ and $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{2d})$. We assume
$\int_{0}^{T}\mathcal{D}(\mu_{t}){\,\rm d}t<+\infty\quad\text{and}\quad\int_{0}^{T}\mathcal{A}(\mu_{t},\mathcal{U}_{t}){\,\rm d}t<+\infty.$ (3.8)
  • Let $\gamma\in(-\gamma_{d},0)$. Let $\mu_{t}=f_{t}\mathcal{L}$ and
    $\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty,\quad\sup_{t\in[0,T]}\mathcal{E}_{0,2+|\gamma|}(\mu_{t})<+\infty,$
    $\text{and}\quad\langle v\rangle^{2+|\gamma|}f\in L^{\infty}([0,T];{L^{p}_{v}L^{1}_{x}})$
    for some $p>\frac{d}{d-|\gamma|}$.
  • Let $\gamma=0$. Let
    $\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty\quad\text{and}\quad\sup_{t\in[0,T]}\mathcal{E}_{0,2}(\mu_{t})<+\infty.$
  • Let $\gamma\in(0,1]$. Let $\mu_{t}=f_{t}\mathcal{L}$ and
    $\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty,\quad\sup_{t\in[0,T]}\mathcal{E}_{0,2+\gamma}(\mu_{t})<+\infty,$
    $\text{and}\quad\langle v\rangle^{\gamma}f\in L^{\infty}([0,T];{L^{p}_{v}L^{1}_{x}})$
    for some $p>\frac{d}{d-\gamma}$.
Then $|\mathcal{H}(\mu_{t})|<+\infty$ and the following chain rule holds
$\mathcal{H}(\mu_{t})-\mathcal{H}(\mu_{s})=\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log f_{r}{\,\rm d}\mathcal{U}_{r}{\,\rm d}r$ (3.9)
for all $0\leq s\leq t\leq T$. Moreover, the map $t\mapsto\mathcal{H}(\mu_{t})$ is absolutely continuous and we have
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log f_{t}{\,\rm d}\mathcal{U}_{t}$ (3.10)
for almost every $t\in[0,T]$.
Remark 3.7.
  • On the domain $\mathbb{T}^{d}\times\mathbb{R}^{d}$, we do not need to assume the moment bounds in $x$, since $\mathbb{T}^{d}$ can be treated as a bounded domain.
  • The assumption (3.8), Lemma 3.3, and the definition of action (3.3) imply that $(\mu_{t},\mathcal{U}_{t})$ are absolutely continuous with densities $(f_{t},U_{t})$. Moreover, the set $\{ff_{*}=0\}$ is negligible for $U_{t}$ for almost every $t$. Hence, the right-hand side of (3.10) is well-defined.
Remark 3.8 (Assumptions on curves).
  • The curve assumptions in Assumption 2.7 imply the curve assumptions in Theorem 3.6. More precisely, in the hard potential case $\gamma\in(0,1]$, in Assumption 2.7, we only assume moment and $L^{p}$-bounds, which in turn ensure the weighted $L^{p}$-bounds required in Theorem 3.6. Let
    $f\in L^{\infty}([0,T];L^{1}_{0,2+\gamma}(\mathbb{R}^{2d}))\cap L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$
    for some $\frac{d}{d-\gamma}<p$, then by Hölder inequality, we have
    $\|\langle v\rangle^{|\gamma|}f\|_{L^{p^{\prime}}_{v}L^{1}_{x}}\leq\|f\|_{L^{p}_{v}L^{1}_{x}}^{\frac{p-p^{\prime}}{p}}\|f\|_{L^{1}_{0,2+|\gamma|}(\mathbb{R}^{2d})}^{\frac{p^{\prime}}{p}},$
    and hence,
    $\langle v\rangle^{\gamma}f\in L^{\infty}([0,T];L^{p^{\prime}}_{v}L^{1}_{x})$
    for all $\max\big(\frac{d}{d-\gamma},\frac{p\gamma}{2+\gamma}\big)<p^{\prime}<p$.
  • There are $L^{p}_{v}L^{1}_{x}$ assumptions on curves in both the soft and hard potential cases, though they play different roles. In the hard potential case $\gamma\in(0,1]$, the $L^{p}$-assumptions are used to bound the weighted Fisher information, see Lemma 2.9. In the soft potential case $\gamma\in(-\gamma_{d},0)$, the $L^{p}$-assumptions are used to control $\|\langle v\rangle^{2}f*_{v}|v|^{\gamma}\|_{L^{\infty}_{v}L^{1}_{x}}$, see, for example, (3.45).
  • Theorem 3.6 holds, if we replace $L^{p}_{v}L^{1}_{x}$ in the assumptions by $L^{1}_{x}L^{p}_{v}$. The $L^{p}$ bound is used to ensure the integrability of $\int_{\mathbb{R}^{4d}}ff_{*}|v-v_{*}|^{\gamma}{\,\rm d}\eta$, for which we use Lemma 2.10, and we make no distinction between $L^{1}_{x}L^{p}_{v}$ and $L^{p}_{v}L^{1}_{x}$ there.
We prove the Theorem 3.6 in three steps: We regularise (in $x$, $v$ and $t$) curves $(\mu_{t},\mathcal{U}_{t})\in\mathcal{TGRE}_{T}$ and discuss the properties of regularised curves in Section 3.2.1. We derive a regularised chain rule and pass to the limit in Section 3.2.2 and Section 3.2.3 respectively.
A similar chain rule for spatial-homogeneous grazing rate equations
$\partial_{t}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0,\quad\mu\in\mathcal{P}(\mathbb{R}^{d}),\quad\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$
has been established in [CDDW24]. In the proof of Theorem 3.6, we follow [CDDW24] to treat the collision terms, where we do appropriate modification concerning the spatial variable $x\in\mathbb{R}^{d}$. We follow [EH25] to treat the transport part $v\cdot\nabla_{x}\mu_{t}$ that appears in the transport grazing rate equations (3.1). The second author and Erbar [EH25] consider a chain rule for a transport collision rate equation related to a delocalised fuzzy Boltzmann equation.
3.2.1 Regularisation
  1. 1.
    Regularisation in $x$.
    Let $M(x)=c\exp\big(-\langle x\rangle\big)$, where we choose the constant $c>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{d})}=1$.
    We define the renormalised sequence of parameter $\alpha\in(0,1)$
    $M_{\alpha}(x)=c\alpha^{-d}\exp\big(-\langle\frac{x}{\alpha}\rangle\big).$
    We define the renormalised $(\mu_{t},\mathcal{U}_{t})$ is the following way
    $\mu^{\alpha}_{t}=\mu_{t}*_{x}M_{\alpha},\quad\mathcal{U}^{\alpha}_{t}=\mathcal{U}_{t}*_{(x,x_{*})}M_{\alpha}.$
    For test functions $\phi:\mathbb{R}^{2d}\to\mathbb{R}$ and $\Phi:\mathbb{R}^{4d}\to\mathbb{R}$, we have
    $\int_{\mathbb{R}^{2d}}\phi{\,\rm d}\mu^{\alpha}_{t}(x,v)=\int_{\mathbb{R}^{3d}}\phi(x-y,v)M_{\alpha}(y){\,\rm d}\mu_{t}(x,v){\,\rm d}y,$
    $\int_{\mathbb{R}^{4d}}\Phi{\,\rm d}\mathcal{U}^{\alpha}_{t}(x,x_{*},v,v_{*})$
    $=\int_{\mathbb{R}^{6d}}\Phi(x-y,x_{*}-y_{*},v,v_{*})M_{\alpha}(y)M_{\alpha}(y_{*}){\,\rm d}\mathcal{U}_{t}(x,x_{*},v,v_{*}){\,\rm d}y{\,\rm d}y_{*}.$
  2. 2.
    Regularisation in $v$.
    We regularise $(\mu^{\alpha}_{t},\mathcal{U}^{\alpha}_{t})$ in $v$ and $v,v_{*}$ by convoluting with $M_{\beta}(v)$, $\beta>0$, in the following way
    $\mu^{\alpha,\beta}_{t}=\mu^{\alpha}_{t}*_{v}M_{\beta},\quad\mathcal{U}^{\alpha,\beta}_{t}=\mathcal{U}^{\alpha}_{t}*_{(v,v_{*})}M_{\beta}.$
  3. 3.
    Regularisation in $t$.
    Let $\eta\in C^{\infty}_{c}(\mathbb{R};\mathbb{R}_{+})$ such that $\operatorname{supp}(\eta)\subset[-1,1]$ and $\int_{\mathbb{R}}\eta=1$. We define the sequence of mollifiers $\eta^{\delta}=\delta^{-1}\eta(\frac{t}{\delta})$. We regularise $(\mu^{\alpha,\beta}_{t},\mathcal{U}^{\alpha,\beta}_{t})$ in time in the following way
    $\mu^{\alpha,\beta,\delta}_{t}=\int_{-\delta}^{\delta}\mu^{\alpha,\beta}_{t-s}\eta^{\delta}_{s}{\,\rm d}s\quad\text{and}\quad\mathcal{U}^{\alpha,\beta,\delta}_{t}=\int_{-\delta}^{\delta}\mathcal{U}^{\alpha,\beta}_{t-s}\eta^{\delta}_{s}{\,\rm d}s$
    for $t\in I^{\delta}\mathop{=}\limits^{\textrm{def}}[\delta,T-\delta]$.
  4. 4.
    A lower bound. We define
    $g(t,x,v)=Z\big(\langle v\rangle+\langle x-tv\rangle\big)^{-a},$ (3.11)
    where we choose $a>8+|\gamma|$ to ensure the integral bounds in Lemma 3.10 hold also for $g$, and we choose $Z>0$ such that $\|g\|_{L^{1}(\mathbb{R}^{2d})}=1$.
    Notice that $g$ solves the transport equation
    $(\partial_{t}+v\cdot\nabla_{x})g=0.$
    There is a positive lower bound
    $g(t,x,v)\geq C(T)\big(\langle v\rangle+\langle x\rangle\big)^{-a}.$
    For $\theta\in(0,1)$, we regularise $(\mu^{\alpha,\beta,\delta}_{t},\mathcal{U}^{\alpha,\beta,\delta}_{t})$ in the following way
    $\mu^{\alpha,\beta,\delta,\theta}_{t}=(1-\theta)\mu^{\alpha,\beta,\delta}_{t}+\theta g\mathcal{L},\quad\mathcal{U}^{\alpha,\beta,\delta,\theta}_{t}=(1-\theta)\mathcal{U}^{\alpha,\beta,\delta}_{t}.$ (3.12)
    We denote $f^{\alpha,\beta,\delta,\theta}_{t}$ and $U^{\alpha,\beta,\delta,\theta}_{t}$ the density of $\mu^{\alpha,\beta,\delta,\theta}_{t}$ and $\mathcal{U}^{\alpha,\beta,\delta,\theta}_{t}$ with respect to Lebesgue measure $\mathcal{L}$ on $\mathbb{R}^{2d}$ and $\mathbb{R}^{4d}$.
Remark 3.9.
  • We use the lower bound regularisation (3.11) to provide a uniform-in-time positive lower bound for the regularised curve, in particular in $x$, since we only time-integrable moment bounds in $x$
    $\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty.$
    The uniform lower bounds are needed, for example, in Lemma 3.10-3.
We have the following properties on the regularised curves $(f^{\alpha,\beta,\delta,\theta}_{t},U^{\alpha,\beta,\delta,\theta}_{t})$.
Lemma 3.10.
  1. 1.
    Mass is preserved
    $\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=\int_{\mathbb{R}^{2d}}f_{t}{\,\rm d}x{\,\rm d}v=1$
    for all $t\in I^{\delta}$.
  2. 2.
    For any fixed $\alpha,\,\beta,\,\delta\geq 0$, we have
    $\int_{I^{\delta}}\mathcal{E}_{a,b}(f^{\alpha,\beta,\delta}_{t}){\,\rm d}t$ $\leq\int_{0}^{T}\mathcal{E}_{a,b}(f_{t}){\,\rm d}t,$ (3.13)
    $\sup_{t\in I^{\delta}}\mathcal{E}_{a,b}(f^{\alpha,\beta,\delta}_{t})$ $\leq\sup_{t\in[0,T]}\mathcal{E}_{a,b}(f_{t}),$ (3.14)
    $\sup_{t\in I^{\delta}}\|f^{\alpha,\beta,\delta}_{t}\|_{L^{q}_{v}L^{1}_{x}}$ $\leq C(T,q)\sup_{t\in[0,T]}\|f_{t}\|_{L^{q}_{v}L^{1}_{x}}$ (3.15)
    for all $a\in[0,2]$, $b\in[0,2+|\gamma|]$ and $q\in[1,p]$.
    For $\theta\geq 0$, the above bounds hold for $(f^{\alpha,\beta,\delta,\theta}_{t},U^{\alpha,\beta,\delta,\theta}_{t})$ up to a constant.
  3. 3.
    Let $\alpha,\,\beta>0$ and $\delta\geq 0$. If $\theta>0$, we have
    $|\log f^{\alpha,\beta,\delta}_{t}(x,v)|$ $\leq C(\theta)\big(1+\log(\langle{v}\rangle+\langle{x}\rangle)\big).$ (3.16)
    If $\theta=0$, we have
    $|\log f^{\alpha,\beta,\delta}_{t}(x,v)|$ $\leq C\big(\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\big)$ (3.17)
    $\text{and}\quad\Big|\mathcal{H}(f^{\alpha,\beta,\delta}_{t})\Big|$ $\leq C\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\mathcal{E}_{1,1}(f^{\delta}_{t})<+\infty$ (3.18)
    for all $(t,x,v)\in I^{\delta}\times\mathbb{R}^{2d}$, where the constant $C=C(\alpha,\beta)>0$.
Proof.
We recall (2.11) that the following pointwise bounds hold
$\big(\langle\cdot\rangle^{p}*M_{\beta}\big)(z)\leq C\langle z\rangle^{p}\quad\forall\,p\in\mathbb{R},$ (3.19)
where $C>0$ is independent of $\beta$.
  1. 1.
    By definition.
  2. 2.
    By definition, we have
    $\int_{I^{\delta}}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $=\int_{I^{\delta}}\int_{\mathbb{R}^{2d}}f^{\delta}_{t}(y,u)\big(\int_{\mathbb{R}^{d}}M_{\alpha}(x-y){\,\rm d}x\big)\big(\int_{\mathbb{R}^{d}}M_{\beta}(v-u)\langle v\rangle^{a}{\,\rm d}v\big){\,\rm d}y{\,\rm d}u{\,\rm d}t$
    $\leq C(\beta,a)\int_{0}^{T}\int_{-\delta}^{\delta}\eta^{\delta}_{s}\int_{\mathbb{R}^{2d}}f_{t-s}(y,u)\langle u\rangle^{a}{\,\rm d}y{\,\rm d}u{\,\rm d}s{\,\rm d}t$
    $\leq C\int_{0}^{T}\int_{\mathbb{R}^{2d}}f_{t}(x,v)\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
    where we use (3.19). The proof for $\langle x\rangle^{b}$ and (3.14) follow analogously.
    We show (3.15). We denote $F^{\beta,\delta}_{t}(v)=\int_{\mathbb{R}^{d}}f^{\beta,\delta}_{t}(x,v){\,\rm d}x$, then we have
    $\|f^{\alpha,\beta,\delta}_{t}\|_{L^{q}_{v}L^{1}_{x}}^{q}$ $=\int_{\mathbb{R}^{d}}\big(F^{\beta,\delta}(v)\big)^{q}{\,\rm d}v$
    $=\int_{\mathbb{R}^{d}}\Big(\int_{-\delta}^{\delta}\int_{\mathbb{R}^{d}}F_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s\Big)^{q}{\,\rm d}v$
    $\leq\int_{\mathbb{R}^{d}}\int_{-\delta}^{\delta}\int_{\mathbb{R}^{d}}F^{q}_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s{\,\rm d}v$
    $\leq\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}F^{q}_{t}(v){\,\rm d}v=\sup_{t\in[0,T]}\|f_{t}\|_{L^{q}_{v}L^{1}_{x}},$
    where we use Jensen’s inequality since $x\mapsto x^{q}$ is convex.
    We only left to show the boundedness of $\mathcal{E}_{a,b}(g)$ and $\|g\|_{L^{q}_{v}L^{1}_{x}}$ ($g$ defined as in (3.11)). Let $h(x,v)=Z\big(\langle v\rangle+\langle x\rangle\big)^{-a}$. By definition (3.11), we have
    $\mathcal{E}_{a,b}(g)\lesssim(1+T)\mathcal{E}_{\max(a,b),b}(h)\quad\text{and}\quad\|g\|_{L^{q}_{v}L^{1}_{x}}=\|h\|_{L^{q}_{v}L^{1}_{x}}.$
  3. 3.
    For fixed $\alpha,\,\beta>0$, the following upper bounds hold
    $f^{\alpha,\beta,\delta}_{t}(x,v)$ $\leq\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.20)
    $\leq(\alpha\beta)^{-1}\int_{\mathbb{R}^{2d}}f^{\delta}{\,\rm d}x{\,\rm d}v=(\alpha\beta)^{-1}.$
    If $\theta>0$, by definition, the following lower bounds hold
    $f^{\alpha,\beta,\delta,\theta}_{t}(x,v)\geq\theta g(t,x,v)\geq C(\theta,T)\big(\langle x\rangle+\langle v\rangle\big)^{-a}.$
    We conclude the bounds (3.16) when $\theta>0$.
    In the case of $\theta=0$, we show the following lower bounds
    $f^{\alpha,\beta,\delta}_{t}(x,v)\geq C\exp\big(-C\big(\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\big),$ (3.21)
    where $C=C(\alpha,\beta)>0$.
    We follow [CDDW24, Lemma 30] with an appropriate modification for the spatial variable $x\in\mathbb{R}^{d}$. We note that
    $\langle\frac{v-w}{\beta}\rangle$ $=\sqrt{1+\big|\frac{v-w}{\beta}\big|^{2}}\leq\sqrt{1+2\big|\frac{v}{\beta}\big|^{2}+2\big|\frac{w}{\beta}\big|^{2}}\leq\sqrt{2}\big(\langle\frac{v}{\beta}\rangle+\langle\frac{w}{\beta}\rangle\big),$
    and similarly
    $\langle\frac{x-y}{\alpha}\rangle\leq\sqrt{2}\big(\langle\frac{x}{\alpha}\rangle+\langle\frac{y}{\alpha}\rangle\big).$
    We define $\Omega_{R}=\{|v|\leq R_{1}\}\times\{|x|\leq R_{2}\}$. We have
    $f^{\alpha,\beta,\delta}(x,v)=\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.22)
    $\geq C(\alpha,\beta)\exp\Big(-\sqrt{2}\big(\langle\frac{v}{\beta}\rangle+\langle\frac{x}{\alpha}\rangle\big)\Big)$
    $\quad\times\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)\exp\Big(-\sqrt{2}\big(\langle\frac{w}{\beta}\rangle+\langle\frac{y}{\alpha}\rangle\big)\Big){\,\rm d}y{\,\rm d}w$
    $\geq C\exp\Big(-\sqrt{2}\big(\langle\frac{v}{\beta}\rangle+\langle\frac{x}{\alpha}\rangle+\langle\frac{R_{1}}{\beta}\rangle+\langle\frac{R_{2}}{\alpha}\rangle\big)\Big)\int_{\Omega_{R}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v.$
    Notice that, on the complement of $\Omega_{R}$, we have
    $\int_{\Omega_{R}^{c}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v$
    $\leq{}$ $\int_{\{|v|\geq R_{1}\}\times\mathbb{R}^{d}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v+\int_{\{|x|\geq R_{2}\}\times\mathbb{R}^{d}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v$
    $\leq{}$ $\int_{\{|v|\geq R_{1}\}\times\mathbb{R}^{d}}\frac{|v|^{2}}{R_{1}^{2}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v+\int_{\{|x|\geq R_{2}\}\times\mathbb{R}^{d}}\frac{|x|^{2}}{R_{2}^{2}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v$
    $\leq{}$ $\frac{\mathcal{E}_{0,2}(f^{\delta}_{t})}{R_{1}^{2}}+\frac{\mathcal{E}_{2,0}(f^{\delta}_{t})}{R_{2}^{2}}.$
    By assumption, we have
    $\sup_{t\in[0,T]}\mathcal{E}_{0,2}(f^{\delta}_{t})<+\infty\quad\text{and}\quad\mathcal{E}_{2,0}(f^{\delta}_{t})<+\infty\quad\forall\,t\in I^{\delta},$
    since $f_{t}\in\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{2d})$ for all $t\in[0,T]$. Then there exist a time-independent $R_{1}$ and time-dependent $R_{2}(t)$ large, such that
    $\frac{\sup_{t}\mathcal{E}_{0,2}(f^{\delta}_{t})}{R_{1}^{2}}\leq\frac{1}{4}\quad\text{and}\quad\frac{\mathcal{E}_{2,0}(f^{\delta}_{t})}{R_{2}^{2}(t)}=\frac{1}{4}\quad\forall\,t\in I^{\delta}.$
    Then we have the following time-dependent lower bound from (3.22)
    $f^{\alpha,\beta,\delta}_{t}(x,v)$ $\geq C(\alpha,\beta)\exp\Big(-\sqrt{2}\big(\langle\frac{v}{\beta}\rangle+\langle\frac{x}{\alpha}\rangle+\langle\frac{R_{1}}{\beta}\rangle+\langle\frac{2\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}}{\alpha}\rangle\big)\Big)$
    $\geq C\exp\big(-C\big(\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\big).$
    The the bound (3.17) is a direct consequence of the upper bound (3.20) and the lower bound (3.21).
    The bounded entropy (3.18) is a direct consequence of (3.17)
    $\Big|\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\log f^{\alpha,\beta,\delta}_{t}\Big|$ $\leq C\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\big(\langle x\rangle+\langle v\rangle\big){\,\rm d}x{\,\rm d}v$
    $\leq C\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\mathcal{E}_{1,1}(f^{\delta}_{t})<+\infty$
    where we use the moment bounds (3.13).
We recall the definition of the entropy dissipation (1.10)
$\mathcal{D}(f)=\frac{1}{2}\int_{\mathbb{R}^{4d}}A\kappa ff_{*}\big|\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)\big|^{2}{\,\rm d}\eta.$ (3.23)
We define $\overline{\nabla}\mathop{=}\limits^{\textrm{def}}\nabla_{v}-\nabla_{v_{*}}$. The following identity holds
$ff_{*}\big|\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)\big|^{2}$
$={}$ $4\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}\sqrt{ff_{*}}\big|^{2}=4\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}.$
Then the entropy dissipation (3.23) can be written as
$\mathcal{D}(f)=2\int_{\mathbb{R}^{4d}}A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}{\,\rm d}\eta.$ (3.24)
We have the following estimates on regularised entropy dissipation and curve action. Notice that we only have the bounds on $\mathcal{D}(f^{\alpha,\delta})$ (with vanishing $\beta=0$), besides in the case in Section 4.1.
Lemma 3.11.
For any fixed $\alpha\geq 0$, we have the pointwise bounds
$A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(f^{\alpha}f^{\alpha}_{*})\big|^{2}}{f^{\alpha}f^{\alpha}_{*}}\leq A\kappa\Big(\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}\Big)*_{x,x_{*}}M_{\alpha}$ (3.25)
and
$\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t\leq C\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t$ (3.26)
For any fixed $\alpha,\,\beta\geq 0$, we have the pointwise bounds
$\frac{|U^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}f^{\alpha,\beta}_{*}}\leq\kappa^{-1}\Big(\frac{|U_{t}|^{2}}{ff_{*}}\Big)*_{v,v_{*}}M_{\beta}*_{x,x_{*}}M_{\alpha}$ (3.27)
and
$\int_{0}^{T}\mathcal{A}(f^{\alpha,\beta,\delta}_{t},U^{\alpha,\beta}_{t}){\,\rm d}t\leq C\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t.$
Proof.
Tow show (3.26), we use the entropy dissipation formulation (3.24)
$\mathcal{D}(f^{\alpha})=2\int_{\mathbb{R}^{4d}}A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(f^{\alpha}f^{\alpha}_{*})\big|^{2}}{f^{\alpha}f^{\alpha}_{*}}{\,\rm d}\eta.$
Notice that the regularisations and $\Pi_{(v-v_{*})^{\perp}}$ and $\overline{\nabla}$ are commuted. By using of the convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$ and Jensen’s inequality we have the pointwise bounds (3.25). By definition of the spatial kernel $\kappa$ in Assumption 2.1 and in Remark 2.2 (we refer to property (2.10) and definition (2.4)), we have
$M_{\alpha}*_{x,x_{*}}\kappa(x-x_{*})\lesssim\kappa(x-x_{*}),$
and
$\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t$ $\leq\int_{0}^{T}\int_{\mathbb{R}^{4d}}A\big(M_{\alpha}*_{x,x_{*}}\kappa\big)\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}{\,\rm d}\eta{\,\rm d}t$
$\leq C\int_{0}^{T}\mathcal{D}(f){\,\rm d}t.$
Similarly, concerning the curve action, we have the pointwise bounds (3.27) and
$\int_{0}^{T}\mathcal{A}(f^{\alpha,\beta},U^{\alpha,\beta}){\,\rm d}t$
$\leq{}$ $\int_{0}^{T}\int_{\mathbb{R}^{4d}}M_{\beta}*_{v,v_{*}}\big(M_{\alpha}*_{x,x_{*}}\kappa^{-1}\big)\frac{|U_{t}|^{2}}{ff_{*}}{\,\rm d}\eta{\,\rm d}t$
$\leq{}$ $C\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{\kappa ff_{*}}{\,\rm d}\eta{\,\rm d}t=C\int_{0}^{T}\mathcal{A}(f,U){\,\rm d}t.$
Corollary 3.12.
For any fixed $\alpha,\,\beta,\,\delta,\,\theta\geq 0$ and $\gamma\in(-\gamma_{d},1]$, we have
$\int_{I^{\delta}}\int_{\mathbb{R}^{4d}}|U_{t}^{\alpha,\beta,\delta,\theta}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big){\,\rm d}\eta{\,\rm d}t<+\infty,$ (3.28)
and
$\int_{I^{\delta}}\int_{\mathbb{R}^{4d}}|U_{t}^{\alpha,\beta,\delta,\theta}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t<+\infty.$ (3.29)
Proof.
In Lemma 3.5, we showed (3.28) and (3.29) for a general grazing rate $U_{t}$. The same bounds hold for regularised $U^{\alpha,\beta,\delta,\theta}_{t}$ by repeating the proof for Lemma 3.5 and using Lemma 3.10 and Lemma 3.11. ∎
3.2.2 Approximated chain rule
The pair $(\mu^{\alpha,\beta,\delta,\theta},\mathcal{U}^{\alpha,\beta,\delta,\theta}_{t})$ is smooth in $t,x,v$ and satisfies the following equation
$(\partial_{t}+v\cdot\nabla_{x})\mu^{\alpha,\beta,\delta,\theta}_{t}+\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}+\frac{1}{2}(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta,\theta})^{\beta}=0,$ (3.30)
where $(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta})^{\beta}=(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta}$. The error term $\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}$ is given by
$\mathsf{E}^{\alpha,\beta,\delta}_{t}=(1-\theta)\int_{\mathbb{R}^{d}}M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u.$
Indeed, we verify (3.30) in the weak formulation with test function $\phi\in C^{\infty}_{c}(\mathbb{R}^{2d})$:
$\frac{d}{dt}\int_{\mathbb{R}^{2d}}\phi{\,\rm d}\mu_{t}^{\alpha,\beta,\delta,\theta}=(1-\theta)\frac{d}{dt}\int_{\mathbb{R}^{2d}}\phi^{\alpha,\beta}{\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi\partial_{t}g{\,\rm d}x{\,\rm d}v,$
and
$-\int_{\mathbb{R}^{2d}}v\cdot\nabla_{x}\phi{\,\rm d}\mu^{\alpha,\beta,\delta,\theta}_{t}$
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha})*_{v}M_{\beta}{\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha}*_{v}M_{\beta}){\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
$\quad-(1-\theta)\int_{\mathbb{R}^{3\times 3}}u\cdot\nabla_{x}(\phi*_{x}M_{\alpha})(v+u)\cdot\nabla M_{\beta}(u){\,\rm d}\mu^{\delta}_{t}(x,v){\,\rm d}u$
$={}$ $-(1-\theta)\int_{\mathbb{R}^{2d}}(v\cdot\nabla_{x}\phi*_{x}M_{\alpha}*_{v}M_{\beta}){\,\rm d}\mu^{\delta}_{t}+\theta\int_{\mathbb{R}^{2d}}\phi(v\cdot\nabla_{x}g){\,\rm d}x{\,\rm d}v$
$\quad+(1-\theta)\int_{\mathbb{R}^{3d}}\phi(v)M_{\beta}(u)u\cdot\nabla_{x}f^{\alpha,\delta}_{t}(x,v-u){\,\rm d}u{\,\rm d}x{\,\rm d}v,$
and finally
$\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\phi^{\alpha,\beta}{\,\rm d}\mathcal{U}^{\delta}_{t}$ $=\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\phi^{\beta}{\,\rm d}\mathcal{U}^{\alpha,\delta}_{t}=-\int_{\mathbb{R}^{2d}}\phi(\widetilde{\nabla}\cdot\mathcal{U}^{\alpha,\delta}_{t})*_{(v,v_{*})}M_{\beta},$
where we use $\Pi_{(v-v_{*})^{\perp}},M_{\alpha}*_{(x,x_{*})}]=0$ and $\cdot$$\cdot$] denotes commutator.
Since $(\mu_{t},\mathcal{U}_{t})\in\mathcal{TGRE}_{T}$ and $(\partial_{t}+v\cdot\nabla_{x})g=0$, we conclude that $(\mu^{\alpha,\beta,\delta}_{t},\mathcal{U}^{\alpha,\beta,\delta}_{t})$ is a pair of solutions to (3.30) at least in the distribution sense.
We verify the approximation chain rule of equation (3.30). We first show the absolute continuity of $t\mapsto\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})$. For any $h>0$ we have
$\frac{1}{h}|f^{\alpha,\beta,\delta,\theta}_{t+h}\log f^{\alpha,\beta,\delta,\theta}_{t+h}-f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}|$
$\leq{}$ $\frac{1}{h}\max\{|\log f^{\alpha,\beta,\delta,\theta}_{t+h}|,|\log f^{\alpha,\beta,\delta,\theta}_{t}|\}|f^{\alpha,\beta,\delta,\theta}_{t+h}-f^{\alpha,\beta,\delta,\theta}_{t}|$
$\leq{}$ $C(\theta)\big(\langle x\rangle+\langle v\rangle\big)\int_{\mathbb{R}}f^{\alpha,\beta}_{r}\Big|\frac{\eta^{\delta}(t+h-r)-\eta^{\delta}(t-r)}{h}\Big|{\,\rm d}r$
$\leq{}$ $C\big(\langle x\rangle+\langle v\rangle\big)\|\eta^{\delta}\|_{\operatorname{Lip}(\mathbb{R})}\int_{0}^{T}f^{\alpha,\beta}_{t}{\,\rm d}t,$
where the bounds on $|\log f^{\alpha,\beta,\delta,\theta}|$ are given in (3.16). The right-hand side of the above inequality is integrable on $\mathbb{R}^{2d}$ by (3.13). We use the dominated convergence theorem to exchange the time derivative and the integral over $\mathbb{R}^{2d}$ to derive
$\frac{d}{dr}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v$
$={}$ $\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta,\delta,\theta}_{r}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{2d}}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v$
$={}$ $\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta,\delta,\theta}_{r}\partial_{r}f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v.$
Hence, we differentiate $\mathcal{H}(f^{\alpha,\beta,\delta}_{t})$ in time to derive the approximated chain rule for all $t\in I^{\delta}$
$\frac{d}{dt}\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})$ $=-\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v$ (3.31)
$\quad+\frac{1}{2}\int_{\mathbb{R}^{4d}}U^{\alpha,\delta,\theta}_{t}\cdot\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{t})^{\beta}{\,\rm d}\eta.$
since $\int_{\mathbb{R}^{2d}}v\cdot\nabla_{x}f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=0$.
We integrate (3.31) in time from $s$ to $t$ to derive
$\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})-\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{s})=$ $-\int_{s}^{t}\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ (3.32)
$+\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{4d}}U^{\alpha,\delta,\theta}_{r}\cdot\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta}_{r})^{\beta}{\,\rm d}\eta{\,\rm d}r$
for all $\delta\leq s\leq t\leq T-\delta$.
3.2.3 Passing to the limit.
We will pass to the limit in order $\delta,\,\theta,\,\beta,\,\alpha\to 0$.
We first recall the dominated convergence theorem here, which will be used frequently below.
Theorem 3.13 (Dominated convergence theorem).
If a pointwise convergence sequence $f_{n}(x)\to f(x)$, for a.e. $x\in X$ is dominated by an integrable sequence $\{g_{n}\}$ in the sense that $|f_{n}(x)|\leq|g_{n}(x)|$, for all $x\in X$ and $n\geq 0$. If $g_{n}\to g$ and $\lim_{n\to\infty}\int_{X}g_{n}=\int_{X}g$, then $\lim_{n\to\infty}\int_{X}f_{n}=\int_{X}f$.
Concerning the time regularisation limit, we will frequently use the following proposition.
Proposition 3.14.
For any $\alpha,\beta\in(0,1)$, we have
$f^{\alpha,\beta,\delta,\theta}_{t}\to f^{\alpha,\beta,\theta}_{t}\quad\text{uniformly on }(0,T)\times\mathbb{R}^{2d}.$
In addition, if $\theta\in(0,1)$, then
$\log f^{\alpha,\beta,\delta,\theta}_{t}\to\log f^{\alpha,\beta,\theta}_{t}\quad\text{uniformly on }(0,T)\times\Omega,$
where $\Omega\subset\mathbb{R}^{2d}$ bounded.
Proof.
We follow [CDDW24].
For any fixed $(t,x,v)$, we have
$|f^{\alpha,\beta,\delta,\theta}(x,v)-f^{\alpha,\beta,\theta}_{t}(x,v)|$
$={}$ $\Big|\int_{\mathbb{R}}f^{\alpha,\beta}_{t-s}\eta^{\delta}(s){\,\rm d}s-\int_{\mathbb{R}}f^{\alpha,\beta}_{t}\eta^{\delta}(s){\,\rm d}s\Big|$
$\leq{}$ $\sup_{s\in\operatorname{supp}(\eta^{\delta})}\Big|\int_{\mathbb{R}^{2d}}(f_{t-s}-f_{t})(y,w)M_{\alpha}(x-y)M_{\beta}(v-w){\,\rm d}y{\,\rm d}w\Big|\to 0$
as a consequence of the uniformly weakly continuity for $(f_{t})_{t\in[0,T]}$.
Concerning the convergence of $\log f^{\alpha,\beta,\delta,\theta}$, we have
$|{\log f_{t}}^{\alpha,\beta,\delta,\theta}(x,v)-\log f_{t}^{\alpha,\beta,\theta}(x,v)|$
$\leq{}$ $(\theta g)^{-1}\big|f^{\alpha,\beta,\delta,\theta}_{t}-f^{\alpha,\beta,\theta}_{t}\big|$
$\leq{}$ $C(T,\theta)\big(\langle x\rangle+\langle v\rangle\big)^{a}\big|f^{\alpha,\beta,\delta,\theta}_{r}(x,v)-f^{\alpha,\beta,\theta}_{r}(x,v)\big|,$
which vanishes uniformly on $(0,T)\times\Omega$. ∎
Collision term
We first show the following estimates for the regularised curves.
Lemma 3.15.
For any fixed $\beta>0$ and $\delta,\,\theta,\,\alpha\geq 0$, we have
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|\lesssim\beta^{-1}\quad\text{and}\quad|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}}.$ (3.33)
In particular, in the case of $\gamma\in[-2,1]$, we have
$|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta}_{t})^{\beta}|\lesssim_{\beta,\gamma}|v|^{1+\frac{\gamma}{2}}+|v_{*}|^{1+\frac{\gamma}{2}}.$
Proof.
The proof follows [CDDW24, Lemma 31 & 32]. We note that
$\nabla_{v}M_{\beta}(v)=-\langle\frac{v}{\beta}\rangle\frac{v}{\beta^{2}}M_{\beta}(v)\quad\text{and}\quad|\nabla_{v}M_{\beta}(v)|\lesssim\beta^{-1}M_{\beta}(v).$ (3.34)
Then we have
$|\nabla_{v}\log f^{\alpha,\beta,\delta}_{t}|$ $=\Big|\frac{\int_{\mathbb{R}^{d}}f^{\alpha,\delta}(x,w)\nabla_{v}M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big|$
$\lesssim\beta^{-1}\Big|\frac{\int_{\mathbb{R}^{d}}f^{\alpha,\delta}(x,w)M_{\beta}(v-w){\,\rm d}w}{f^{\alpha,\beta,\delta}_{t}}\Big|\lesssim\beta^{-1}.$
Combining with $|\nabla_{v}g|\lesssim 1$ and $\beta\in(0,1)$, we have the first bound in (3.33).
To show the second bound in (3.33), by definition, we have
$\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}=$ $|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}-\nabla_{v_{*}}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}\big).$
We note that
$\big|\nabla_{v}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}\big|=|(\nabla_{v}\log f^{\alpha,\beta,\delta,\theta}_{t})*_{v}M_{\beta}|\lesssim\beta^{-1}\|M_{\beta}\|_{L^{1}}.$
  • Limit $\delta,\,\theta\to 0$.
    We will show
    $\lim_{\theta\to 0}\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}\cdot U_{r}^{\alpha,\delta,\theta}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta})^{\beta}\cdot U_{r}^{\alpha}{\,\rm d}\eta{\,\rm d}r.$
    To show the $\delta$-limit, we only need to show as $\delta\to 0$
    $(1-\theta)\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\cdot\Big(U_{r}^{\alpha,\delta}-U_{r}^{\alpha}\Big){\,\rm d}\eta{\,\rm d}r\Big|\to 0\quad\text{and}$ (3.35)
    $(1-\theta)\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\Big(\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}\Big)\cdot U_{r}^{\alpha,\delta}{\,\rm d}\eta{\,\rm d}r\Big|\to 0.$ (3.36)
    We have the uniformly bound (3.33) and the integrability (3.28), i.e.
    $|\widetilde{\nabla}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}|\lesssim_{\beta}|v-v_{*}|^{1+\frac{\gamma}{2}},\quad|U^{\alpha,\beta,\delta}||v-v_{*}|^{1+\frac{\gamma}{2}}\in L^{1}([0,T]\times\mathbb{R}^{4d}).$ (3.37)
    Hence, by definition (3.35) vanishes and we only need to show the vanishing of (3.36) on a bounded domain $B^{4}_{l}(0)$ for some large $l$.
    We define the ball $B^{m}_{k}(\cdot)\subset\mathbb{R}^{md}$ as the ball centred at $\cdot\in\mathbb{R}^{md}$ with radius $k$.
    To estimate (3.36), we combine $\nabla_{v}$ and $\nabla_{v_{*}}$ terms in the following way
    $|\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\delta,\theta})^{\beta}-\widetilde{\nabla}(\log f_{r}^{\alpha,\beta,\theta})^{\beta}|$
    $\leq|v-v_{*}|^{1+\frac{\gamma}{2}}\big(|\nabla_{v}(\log f^{\alpha,\beta,\delta,\theta})^{\beta}-\nabla_{v}(\log f^{\alpha,\beta,\theta})^{\beta}|$
    $\quad+|\nabla_{v_{*}}(\log f_{*}^{\alpha,\beta,\delta,\theta})^{\beta}-\nabla_{v_{*}}(\log f_{*}^{\alpha,\beta,\theta})^{\beta}|\big).$
    We only need to treat the $\nabla_{v}$ in the integral due to the changing of variables. For $(x,v)\in B^{2}_{l}(0)$, we note that
    $\nabla_{v}\big(\log f^{\alpha,\beta,\delta,\theta}\big)^{\beta}-\nabla_{v}\big(\log f^{\alpha,\beta,\theta}\big)^{\beta}$
    $={}$ $\nabla_{v}\Big(\big(\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big)*_{v}M_{\beta}\Big)$
    $={}$ $\int_{B^{1}_{2l}(0)}\big(\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big)(u)\nabla_{v}M_{\beta}(v-u){\,\rm d}u$
    $+\int_{\big(B^{1}_{2l}(v)\big)^{c}}\nabla_{v}\big(\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big)(v-u)M_{\beta}(u){\,\rm d}u.$
    Notice that $\big(B^{1}_{2l}(v)\big)^{c}\subset\big(B^{1}_{l}(0)\big)^{c}$. As a consequence of the pointwise bounds (3.33) $|\nabla_{v}\log f^{\alpha,\beta,\delta,\theta}|\lesssim\beta^{-1}$ and $\|M_{\beta}\|_{L^{1}}=1$, the integral over $(B^{1}_{2l}(v))^{c}$ vanishes as $l\to\infty$.
    On the bounded domain $\Omega\mathop{=}\limits^{\textrm{def}}B_{l}^{4}(0)\times B^{1}_{2l}(0)$, we have
    $\Big|\int_{s}^{t}\int_{\Omega}|v-v_{*}|^{1+\frac{\gamma}{2}}\big|U^{\alpha,\delta}\big|({\log f^{\alpha,\beta,\delta,\theta}}-\log f^{\alpha,\beta,\theta})(u)\nabla M_{\beta}(v-u){\,\rm d}u{\,\rm d}\eta{\,\rm d}r\Big|$
    $\leq\sup_{(t,x,u)\in I^{\delta}\times B_{2l}^{2}(0)}\big|\log f^{\alpha,\beta,\delta,\theta}-\log f^{\alpha,\beta,\theta}\big|\|\nabla M_{\beta}\|_{L^{\infty}}$
    $\quad\times\||v-v_{*}|^{1+\frac{\gamma}{2}}U^{\alpha,\delta}\|_{L^{1}([0,T]\times\mathbb{R}^{4d})}\to 0\quad\text{as}\quad\delta\to 0,$
    where we use the uniform convergence of $\log f^{\alpha,\beta,\delta}$ given in Proposition 3.14.
    We use the dominated convergence Theorem 3.13 to show the $\theta$-limit. Notice that the bounds (3.37) imply that $\big|\widetilde{\nabla}\big(\log f^{\alpha,\beta,\theta}\big)^{\beta}\cdot U^{\alpha,\theta}\big|\lesssim\beta^{-1}|U^{\alpha}|$.
  • Limit $\beta\to 0$.
    We show that
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r.$ (3.38)
    We follow [CDDW24]. By Cauchy–Schwarz inequality, we have
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big|$
    $\leq\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha}){\,\rm d}t,$
    where the boundedness of the curve action is ensured by Lemma 3.11.
    Similar to (2.16), the product identity $|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=\frac{1}{2}\sum_{i,j=1}^{d}|x_{i}y_{j}-x_{j}y_{i}|^{2}$ implies that
    $\Big|\widetilde{\nabla}(\log f^{\alpha,\beta})^{\beta}\Big|^{2}$
    $=$ ${}\frac{1}{2}\sum_{i,j=1}^{d}|v-v_{*}|^{\gamma}\Big|\Big[v-v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}\Big]_{ij}\Big|^{2}.$
    The cross-product term has the following bounds
    $\Big|\Big[v-v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}-\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}\Big]_{ij}\Big|^{2}$
    $\leq{}$ $4\Big(|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v_{*}}\log f_{*}^{\alpha,\beta}*_{v_{*}}M_{\beta}]_{ij}|^{2}$
    $\quad+|[v,\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}\Big)$
    $\lesssim{}$ $\Big(|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}+|[v_{*},\nabla_{v_{*}}\log f_{*}^{\alpha,\beta}*_{v_{*}}M_{\beta}]_{ij}|^{2}$
    $\quad+|v|^{2}|\nabla_{v_{*}}\log f^{\alpha,\beta}_{*}*_{v_{*}}M_{\beta}|^{2}+|v_{*}|^{2}|\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}|^{2}\Big).$
    As $\beta\to 0$, the upper bound pointwise converges to
    $\frac{|[v,\nabla_{v}f^{\alpha}]_{ij}|^{2}}{(f^{\alpha})^{2}}+\frac{|[v_{*},\nabla_{v_{*}}f_{*}^{\alpha}]_{ij}|^{2}}{(f_{*}^{\alpha})^{2}}+|v|^{2}\frac{|\nabla_{v_{*}}f^{\alpha}_{*}|^{2}}{(f^{\alpha}_{*})^{2}}+|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{(f^{\alpha})^{2}},$ (3.39)
    which refer to the Fisher and the cross-Fisher information in Lemma (2.9).
    To show (3.38), we only need to show the convergence
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.40)
    $=\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}\frac{|[v,\nabla_{v}f^{\alpha}]_{ij}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t$
    and
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}|\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.41)
    $=\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t.$
    Notice that the limits corresponding to the second and third terms in (3.39) follow analogy.
    We first show the convergence (3.40). By using of the identity $v-w,\nabla M_{\beta}(v-w)]_{ij}=0$, we have
    $[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}=\Big[v,\int_{\mathbb{R}^{d}}\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w){\,\rm d}w\Big]$
    $=\int_{\mathbb{R}^{d}}[w,\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w)]_{ij}{\,\rm d}w=\big([v,\nabla_{v}\log f^{\alpha,\beta}]_{ij}\big)*_{v}M_{\beta}.$
    By Cauchy–Schwarz inequality, we have
    $|\big([v,\nabla_{v}\log f^{\alpha,\beta}]_{ij}\big)*_{v}M_{\beta}|^{2}$ (3.42)
    $\leq{}$ $\Big(\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{-\gamma}{\,\rm d}w\Big)$
    $\quad\times\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\big|[w,\nabla_{w}\log f^{\alpha,\beta}]_{ij}\big|^{2}{\,\rm d}w$
    $\lesssim{}$ $\langle v\rangle^{-\gamma}\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\Big|\frac{[w,\nabla_{w}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}{\,\rm d}w,$
    where we use (3.19) for the last inequality.
    Back to the left-hand side of (3.40), we have
    $\int_{0}^{T}\int_{\mathbb{R}^{3\times 4}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}{\,\rm d}\eta{\,\rm d}t$
    $\lesssim{}$ $\int_{0}^{T}\int_{\mathbb{R}^{2d}}\Big(\int_{\mathbb{R}^{2d}}f^{\alpha}_{*}|v-v_{*}|^{\gamma}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big)f^{\alpha}\langle v\rangle^{-\gamma}$
    $\quad\times\Big(\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\Big|\frac{[w,\nabla_{w}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}{\,\rm d}w\Big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\lesssim{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,0}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}f^{\alpha}M_{\beta}*_{v}\Big(\langle v\rangle^{\gamma}\times\Big|\frac{[v,\nabla_{v}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}\Big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $={}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,0}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\times\frac{|[v,\nabla_{v}f^{\alpha,\beta}]_{ij}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
    where we use Lemma 2.10 and Lemma 3.10, and the constant $C^{\pm}_{\gamma,0}$ is given in (2.24) and (2.25).
    Notice that the joint convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$ and Jensen’s inequality imply that
    $\langle v\rangle^{\gamma}\frac{|[v,\nabla_{v}f^{\alpha}*_{v}M_{\beta}|^{2}]_{ij}}{f^{\alpha}*_{v}M_{\beta}}\leq\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}.$ (3.43)
    Notice that $M_{\beta}*_{v}\langle v\rangle^{\gamma}\lesssim\langle v\rangle^{\gamma}$. Hence, we have
    $\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}{\,\rm d}x{\,\rm d}v,$ (3.44)
    where the right-hand side of the above inequality is bounded by entropy dissipation in Lemma 2.9. By dominated convergence theorem 3.13, we conclude with the convergence (3.40).
    The convergence (3.41) holds similarly to (3.40).
    Similar to (3.42), we have
    $|\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}|^{2}\lesssim\langle v\rangle^{-\gamma}M_{\beta}*_{v}\big(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2}\big).$
    Then the left-hand side of (3.41) is bounded by
    $\int_{0}^{T}\int_{\mathbb{R}^{2d}}\Big(\int_{\mathbb{R}^{2d}}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big)$ (3.45)
    $\quad\times f^{\alpha}\langle v\rangle^{-\gamma}\big(M_{\beta}*_{v}(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2})\big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\int_{0}^{T}C^{\pm}_{\gamma,2}(f^{\alpha}_{t})\int_{\mathbb{R}^{2d}}f^{\alpha}M_{\beta}*_{v}\big(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2}\big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,2}(f^{\alpha}_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta}\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,2}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
    where we use Lemma 2.10 and Lemma 3.10, and the constant $C^{\pm}_{\gamma,2}$ is given in (2.24) and (2.25).
    Similar to (3.43)-(3.44), we have the following bounds
    $\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}\leq\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|\nabla_{v}f|^{2}}{f}$
    $\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f|^{2}}{f}{\,\rm d}x{\,\rm d}v.$
    The right-hand side of the second inequality if bounded by $\mathcal{D}(f)$ shown in Lemma 2.9. Hence, the dominated convergence Theorem 3.13 implies the limit (3.41).
  • Limit $\alpha\to 0$. By Cauchy–Schwarz inequality
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big|\leq\frac{1}{2}\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha},U^{\alpha}){\,\rm d}t.$
    To show the limit of $\alpha\to 0$, we use the pointwise bounds (3.25) and (3.27) in Lemma 3.11 and the dominated convergence Theorem 3.13.
Error term
  • Limit $\delta\to 0$.
    We show that
    $\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=(1-\theta)\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
    Similar to (3.34), we have
    $|\nabla_{x}M_{\alpha}(x)|\lesssim\alpha^{-1}M_{\alpha}(x).$
    Combining with the uniform-in-time bounds (3.17) for $|\log f^{\alpha,\beta,\delta,\theta}|$, we have
    $|\log f^{\alpha,\beta,\delta}_{r}u\cdot\nabla_{x}f^{\alpha,\delta}_{r}(v-u)M_{\beta}(u)|$
    $\leq C(\theta,\alpha)\Big(\int_{\mathbb{R}^{d}}\big(\langle v\rangle+\langle x\rangle\big)|u|M_{\beta}(u)M_{\alpha}(y)f^{\delta}(x-y,v-u){\,\rm d}y\Big)$
    Notice that
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle+\langle x\rangle\big)|u|M_{\beta}(u)M_{\alpha}(y)|f^{\delta}-f|(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq\|\langle v\rangle^{2}M_{\beta}\|_{L^{1}}\|\langle x\rangle^{2}M_{\alpha}\|_{L^{1}}\Big|\int_{0}^{T}\int_{\mathbb{R}^{2d}}\big(\langle v\rangle+\langle x\rangle\big)(f^{\delta}_{t}-f_{t}){\,\rm d}x{\,\rm d}v{\,\rm d}t\Big|,$
    which vanishes as $\delta\to 0$, since $f\in L^{1}([0,T];L^{1}_{1,1}(\mathbb{R}^{2d}))$. By the dominated convergence theorem, we obtain the limit as $\delta\to 0$.
  • Limit $\theta\to 0$.
    To show the $\theta$-limit, we only need to show the following integrability
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq$ ${}C(\alpha,\beta)\int_{s}^{t}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f_{t})^{\frac{1}{2}}\big)$
    $\quad\times|u|M_{\beta}(u)M_{\alpha}(y)f(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq$ ${}C(\alpha,\beta,T)\|\langle v\rangle^{2}M_{\beta}\|_{L^{1}}\|\langle x\rangle M_{\alpha}\|_{L^{1}}\int_{0}^{T}\mathcal{E}_{2,1}(f_{t}){\,\rm d}t<+\infty,$
    where we use the pointwise-in-time bounds (3.17) for $|\log f^{\alpha,\beta}_{t}|$ in the first inequality.
  • Limit $\beta\to 0$.
    We will show that
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r=0.$
    The integration by parts in $x$ implies that
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=-\int_{s}^{t}\int_{\mathbb{R}^{3d}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
    By using Cauchy–Schwarz inequality, one has
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot\int_{\mathbb{R}^{d}}uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq\Big(\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big)^{\frac{1}{2}}\times$
    $\quad\times\Big(\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{(\int_{\mathbb{R}^{d}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{d}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big)^{\frac{1}{2}}.$
    By using of the joint convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$ and Jensen’s inequality, one has
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{(\int_{\mathbb{R}^{d}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{d}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq{}$ $\int_{s}^{t}\int_{\mathbb{R}^{3d}}\frac{(|u|f^{\alpha}_{r}(v-u))^{2}}{f^{\alpha}_{r}(v-u)}M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq{}$ $\|f\|_{L^{1}([0,T]\times\mathbb{R}^{2d})}\int_{\mathbb{R}^{d}}{|v|^{2}M_{\beta}(v)}{\,\rm d}v$
    $\leq{}$ $CT\beta^{2}.$
    On the other hand, similarly, by using the joint convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$, we have the uniform bounds
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ $\leq\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha}_{r}|^{2}}{f^{\alpha}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq\sup_{t\in[0,T]}C(t)\Big(T+\int_{0}^{T}\mathcal{D}(f){\,\rm d}t\Big),$
    where we use Lemma 2.9 in the last step.
    We conclude that, for a fixed $\alpha>0$, we have
    $|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r|$
    $\leq$ ${}C(T)\beta\to 0\quad\text{as}\quad\beta\to 0.$
Entropy term
We first treat the limit $\lim_{\delta\to 0}\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})$ for all $t\in(0,T)$.
The convexity of $x\mapsto x\log x$ implies that
$|\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})-\mathcal{H}(f_{t}^{\alpha,\beta,\theta})|$ (3.46)
$\leq{}$ $\int_{\mathbb{R}^{2d}}|f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}-f_{t}^{\alpha,\beta,\theta}\log f_{t}^{\alpha,\beta,\theta}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $\int_{\mathbb{R}^{2d}}|f^{\alpha,\beta,\delta,\theta}_{t}-f_{t}^{\alpha,\beta,\theta}|\sup_{\delta}|\log f^{\alpha,\beta,\delta,\theta}_{t}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C(\theta)\int_{\mathbb{R}^{2d}}\big(\langle v\rangle+\langle x\rangle\big)|f^{\alpha,\beta,\delta}_{t}-f_{t}^{\alpha,\beta}|{\,\rm d}x{\,\rm d}v\to 0$
as $\delta\to 0$, where the uniform bounds of $|\log f^{\alpha,\beta,\delta,\theta}|$ is given by (3.16). Indeed, the uniform integrability of the right-hand side of the above inequality is given by Lemma 3.10-2, and the uniform convergence of $f^{\alpha,\beta,\delta}$ is given in (3.14).
Notice that by definition $\mathcal{H}(f^{\alpha,\beta,\theta}_{t})$ is well-defined for $t\in[0,T]$. To show the $\theta\to 0$ limit, we can repeat the arguments in (3.46), where we use the pointwise-in-time bounds for $|\log f^{\alpha,\beta}|$ for a fixed time $t$ given by (3.17).
We left to show the $\alpha,\,\beta\to 0$ limits. Since $\mathcal{H}(f^{\alpha,\beta}_{t})$ is lower semicontinuous and $f^{\alpha,\beta}_{t}\to f_{t}$ in $L^{1}_{2,2}(\mathbb{R}^{2d})$ as $\alpha,\,\beta\to 0$, we have
$\mathcal{H}(f_{t})\leq\liminf_{\alpha,\beta\to 0}\mathcal{H}(f^{\alpha,\beta}_{t}).$
On the other hand, the convexity of $a\mapsto a\log a$ and Jensen’s inequality imply for any $\alpha,\,\beta\in(0,1)$
$\mathcal{H}\Big(\int_{\mathbb{R}^{2d}}f_{t}(x-y,v-u)M_{\alpha}(y)M_{\beta}(u){\,\rm d}y{\,\rm d}u\Big)\leq\mathcal{H}(f_{t}).$
Hence, we conclude that $\lim_{\alpha,\beta\to 0}\mathcal{H}(f^{\alpha,\beta}_{t})=\mathcal{H}(f_{t})$. We choose $s=0$ in the chain rule (3.9), the boundedness of $\mathcal{H}(\mu_{0})$ and the right-hand side of the chain rule ensures that $\mathcal{H}(f_{t})$ is bounded and $t\mapsto\mathcal{H}(f_{t})$ is absolutely continuous.
3.3 Variational characterisation for the fuzzy Landau equation
We are ready to prove our main theorem.
Theorem 3.16.
Let $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$. Let $(f_{t})_{t\in[0,T]}$ satisfy the Assumption 2.7. Then we have
$\mathcal{J}_{T}(f,U):=\mathcal{H}(f_{T})-\mathcal{H}(f_{0})+\frac{1}{2}\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t\geq 0.$
Furthermore, $\mathcal{J}_{T}(f,U)=0$ if and only if $f$ is a $\mathcal{H}$-solution of the fuzzy Landau equation (1.1).
Proof.
We first observe that $\mathcal{H}(f_{t})>$-$\infty$ for all $t\in[0,T]$. We follow [JKO98], for some $\kappa>0$, Hölder inequality implies
$\int_{\mathbb{R}^{2d}}|{\min\{f\log f,0\}}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C\int_{\mathbb{R}^{2d}}f^{1-\kappa}{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C\Big(\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{2}+\langle v\rangle^{2}\big)^{-\frac{1-\kappa}{\kappa}}{\,\rm d}x{\,\rm d}v\Big)^{\kappa}\Big(\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{2}+\langle v\rangle^{2}\big)f{\,\rm d}x{\,\rm d}v\Big)^{1-\kappa}$
where we choose $\kappa>0$ small such that $(\langle x\rangle^{2}+\langle v\rangle^{2})^{-\frac{1-\kappa}{\kappa}}\in L^{1}(\mathbb{R}^{2d})$. Hence, we have
$\mathcal{H}(f_{t})\geq-C\mathcal{E}_{2,2}(f_{t})^{1-\kappa}\quad\forall\,t\in[0,T].$ (3.47)
We show $\mathcal{J}_{T}(f,U)\geq 0$. As a consequence of (3.47) and the assumption $|\mathcal{H}(f_{0})|<+\infty$, we have $\mathcal{H}(f_{T})-\mathcal{H}(f_{0})>$-$\infty$. On the other hand, if $\mathcal{D}(f_{t})$ and $\mathcal{A}(f_{t})$ are time integrable, then $\mathcal{J}_{T}(f,U)\geq 0$ is a direct consequence of the chain rule Theorem 3.6, since
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t\geq-\frac{1}{2}\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t-\frac{1}{2}\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t.$
We consider the curves such that $J_{T}(f)=0$. Notice that $J_{T}(f)=0$ and $\mathcal{H}(f_{T})-\mathcal{H}(f_{0})>$-$\infty$ imply the boundedness $\int_{0}^{T}\mathcal{D}(f_{t})+\mathcal{A}(f_{t},U_{t}){\,\rm d}t<+\infty$. The chain rule Theorem 3.6 yields that
$\mathcal{H}(f_{T})-\mathcal{H}(f_{0})=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t$
$\geq-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big(-\widetilde{\nabla}\log(f_{t})\big)^{2}ff_{*}\kappa{\,\rm d}\eta{\,\rm d}t-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U_{t}|^{2}}{{ff_{*}\kappa}}{\,\rm d}\eta{\,\rm d}t.$
We note that $J_{T}(f)=0$ implies that the second inequality must hold as an equality, which means that
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big|\widetilde{\nabla}\log(f_{t})\sqrt{ff_{*}\kappa}+\frac{U_{t}}{\sqrt{ff_{*}\kappa}}\big|^{2}{\,\rm d}\eta{\,\rm d}t=0.$
Hence, we have
$U=-ff_{*}\kappa\widetilde{\nabla}\log f$
almost everywhere on $0,T]\times\mathbb{R}^{4d}$. The weak formulation (2.13) for the fuzzy Landau equation holds since $(f,-\kappa ff_{*}\widetilde{\nabla}\log f)\in\mathcal{TGRE}_{T}$.
If $f_{t}$ is a $\mathcal{H}$-solutions of (1.1), (notice that the entropy dissipation $\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t<+\infty$ by definition), we take $U=-\kappa ff_{*}\widetilde{\nabla}\log f$, then $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$ and
$\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t=\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t<+\infty.$
The chain rule in Theorem 3.6 ensures $\mathcal{J}_{T}(f)=0$.
4 Further discussion and outlook
In this section, we provide further discussion on the GENERIC structure of the fuzzy Landau equation and on possible directions for future works.
4.1 Other GENERIC building blocks
It has been known that an evolution equation may possess different GENERIC structures [Mie23]. For instance, the equation $\partial_{t}\rho=\Delta\rho$ can be written as a gradient flow (that is a GENERIC system in the absence of the reversible component) in at least two different ways
$\partial_{t}\rho=\Delta\rho=\mathsf{M}_{1}(\rho){\,\rm d}\mathsf{S}_{1}(\rho)=\mathsf{M}_{2}(\rho){\,\rm d}\mathsf{S}_{2}(\rho),$
where
$\mathsf{M}_{1}(\rho)\xi=-\mathrm{div}(\rho\nabla\xi),~\mathsf{S}_{1}(\rho)=\int\rho\log\rho$
$\text{and}\quad\mathsf{M}_{2}(\rho)\xi=-\mathrm{div}(\rho^{2}\nabla\xi),~\mathsf{S}_{2}(\rho)=\int\log\rho.$
These two gradient flow structures are respectively related to (via large deviation principle) microscopic models of diffusion and heat conduction processes [MPR14]. The Boltzmann equation can also be formulated in two different GENERIC structures as in [Grm18] and [Ött97].
We now formally show that the (fuzzy) Landau equation also have different GENERIC structures. In fact, the definition of the fuzzy Landau gradient (1.11)
$\widetilde{\nabla}f=\sqrt{A(v-v_{*})}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big)$
can be generalised to the following families of gradient
$\widetilde{\nabla}_{\xi}f=\xi(|v-v_{*}|)\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big)$
$\text{and}\quad\widetilde{\nabla}_{\Xi}f=\Xi(v-v_{*})\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big),$
where
$\xi=\xi(|v-v_{*}|):\mathbb{R}_{+}\to\mathbb{R}_{+}\quad\text{and}$
$\Xi=\Xi(v-v_{*}):\mathbb{R}^{d}\to\mathbb{R}^{d\times d}\quad\text{such that}\quad\Xi=\Xi^{T}.$
Moreover, there exists $\xi_{i}=\xi_{i}(v-v_{*}):\mathbb{R}^{d}\to\mathbb{R}_{+}$, $i=1,2$, such that
$\xi^{2}(|v-v_{*}|)\xi_{1}(v-v_{*})=A(v-v_{*}),$
$\xi_{2}(v-v_{*})\Xi^{T}\Pi_{(v-v_{*})^{\perp}}\Xi=A(v-v_{*})\Pi_{(v-v_{*})^{\perp}}.$
Then the corresponding divergence operators, $\widetilde{\nabla}_{\xi}\cdot$ and $\widetilde{\nabla}_{\Xi}\cdot$, are given by
$\widetilde{\nabla}_{\xi}\cdot G=\nabla_{v}\cdot\int_{\mathbb{R}^{2d}}\xi_{1}(|v-v_{*}|)\Pi_{(v-v_{*})^{\perp}}(G-G_{*}){\,\rm d}x_{*}{\,\rm d}v_{*},$
$\widetilde{\nabla}_{\Xi}\cdot G=\nabla_{v}\cdot\int_{\mathbb{R}^{2d}}\Xi(v-v_{*})^{T}(G-G_{*}){\,\rm d}x_{*}{\,\rm d}v_{*},$
for $G:\mathbb{R}^{4d}\to\mathbb{R}^{d}$. Notice that one can also include the spatial kernel $\kappa(x-x_{*})$ into the above gradient operators. For simplicity of notations, we take $\kappa=1$ in the following of this section.
Let $\widetilde{\nabla}_{\xi/\Xi}$ denote the Landau gradient operator $\widetilde{\nabla}_{\xi}$ or $\widetilde{\nabla}_{\Xi}$. The fuzzy Landau equation (1.1) admits infinite many GENERIC building blocks $\{\mathsf{M}_{\xi/\Xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ where $\mathsf{S}$, $\mathsf{E}$ and $\mathsf{L}$ are given as in Section 1.3, and the operator $\mathsf{M}_{\xi/\Xi}$ is replaced by
$\mathsf{M}_{\xi}(f)g=-\frac{1}{2}\widetilde{\nabla}_{\xi}\cdot\big(\xi_{1}ff_{*}\widetilde{\nabla}_{\xi}g\big)\quad\text{and}$
$\mathsf{M}_{\Xi}(f)g=-\frac{1}{2}\widetilde{\nabla}_{\Xi}\cdot\big(\xi_{2}ff_{*}\Pi_{(v-v_{*})^{\perp}}\widetilde{\nabla}_{\Xi}g\big)$
for all $g\in\mathsf{Z}$. The building block associated to $\mathsf{M}_{\Xi}$ has been proposed in [GGPTZ25].
The variational characterisation in Theorem 1.1 holds for building blocks $\{\mathsf{M}_{\xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ and $\{\mathsf{M}_{\Xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ at least on a formal level. To rigorously establish the GENERIC structure, the key step is to prove the chain rule
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}_{\xi/\Xi}\log f_{t}{\,\rm d}\mathcal{U}_{t}.$
In this article, we have shown the chain rule for $\widetilde{\nabla}_{\xi}$ for the case of $\xi=\sqrt{A}$. Other possible approach is to follow the strategy developed for the Boltzmann equation [Erb23, EH25]. This relies on the commutation property $\widetilde{\nabla}_{\xi},M_{\beta}*_{v}]=0$, that is,
$\big(\widetilde{\nabla}_{\xi}\cdot U\big)*_{v}M_{\beta}=\widetilde{\nabla}_{\xi}\cdot\big(U*_{v,v_{*}}M_{\beta}\big).$
However, it is unclear whether such identity holds in general. This identity may hold for $\widetilde{\nabla}_{\Xi}$ in special cases; for instance, one can take $\Xi=\operatorname{Id}$. Nevertheless, in order to control the collision term, we lack of estimates in the direction of $v-v_{*}$.
4.2 Other kinetic models
Together with previous works [CDDW24, Erb23, EH25] on the Boltzmann/Landau equations, our results cast the spatial homogeneous, fuzzy (delocalised), and classical (at least formally) versions of these equations into a common variational framework. This paves the way for deriving one equation from another by passing to the limit a certain parameter (such as from the Boltzmann equations to the Landau equations in the grazing limit, or from the fuzzy models to the classical ones in the regularisation limits) by generalising the well-established evolutionary $\Gamma$-convergence methods for gradient flows [SS04, Mie16] to GENERIC systems. Another interesting direction for future work is to generalise the results of this paper to other kinetic models such as the relativistic/quantum Boltzmann/Landau equations, and their multi-species models.
Appendix A GENERIC formulation
In this appendix, we show how formally fuzzy Landau equation (1.1) can be cast into a GENERIC framework
$\partial_{t}f=\mathsf{L}(f){\,\rm d}\mathsf{E}(f)+\mathsf{M}(f){\,\rm d}\mathsf{S}(f).$ (A.1)
We consider the space $\mathsf{Z}$ to be the space that consists of all Schwartz functions $f\in\mathcal{S}(\mathbb{R}^{2d})$ such that $f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{2d})$. The space is endowed with the $L^{2}$-inner product $\langle f,g\rangle=\int_{\mathbb{R}^{2d}}fg{\,\rm d}x{\,\rm d}v$. We formally consider the tangent and co-tangent space at $f\in\mathsf{Z}$ as $T_{f}\mathsf{Z}=\{\xi\in L^{2}(\mathbb{R}^{2d})\mid\int_{\mathbb{R}^{2d}}\xi{\,\rm d}x{\,\rm d}v=0\}$ and $T_{f}^{*}\mathsf{Z}=L^{2}(\mathbb{R}^{2d})\}$.
In the case of the fuzzy Landau equation (1.1), we define the energy and entropy functionals by
$\mathsf{E}(f)=\int_{\mathbb{R}^{2d}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f).$
We define the operators $\mathsf{L}$ and $\mathsf{M}$ at $f\in\mathsf{Z}$ by
$\mathsf{M}(f)g$ $=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)\quad\text{and}$
$\mathsf{L}(f)g$ $=-\nabla\cdot(f\mathsf{J}\nabla g),\quad\mathsf{J}=\begin{pmatrix}0&\mathsf{id}_{d}\\ -\mathsf{id}_{d}&0\end{pmatrix}$
for all $g\in\mathsf{Z}$, where $\nabla=(\nabla_{x},\nabla_{v})^{T}$.
We can check directly that $\{\mathsf{L},\mathsf{M},\mathsf{E},\mathsf{S}\}$ is a GENERIC building block for the fuzzy Landau equation (1.1). Indeed, $\mathsf{L}$ is anti-symmetric and satisfies the Jacobi identity, see [DPZ13]
$\langle\mathsf{L}(f)g,g\rangle=\int_{\mathbb{R}^{2d}}f\nabla g\mathsf{J}\nabla g=0$
for all $g\in\mathsf{Z}$. The operator $\mathsf{M}$ is symmetric, positive semi-definite since, using the integration by parts formula (1.12), we have
$\langle\mathsf{M}(f)g,h\rangle$ $=-\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)h{\,\rm d}\eta$
$=\frac{1}{2}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}g\cdot\widetilde{\nabla}h{\,\rm d}\eta=\langle\mathsf{M}(f)h,g\rangle$
$\text{and}\quad\langle\mathsf{M}(f)g,g\rangle$ $=\frac{1}{2}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}g|^{2}{\,\rm d}\eta\geq 0$
for all $g,h\in\mathsf{Z}$.
We indeed recover the fuzzy Landau equation (1.1) from the GENERIC equation (A.1) since
  • ${\,\rm d}\mathsf{E}(f)=\frac{|v|^{2}}{2}$ and
    $\mathsf{L}(f){\,\rm d}\mathsf{E}(f)$ $=-\frac{1}{2}\nabla\cdot\big(f\mathsf{J}\nabla|v|^{2}\big)=-v\cdot\nabla_{x}f,$
  • ${\,\rm d}\mathsf{S}(f)=-(\log f+1)$ and
    $\mathsf{M}(f){\,\rm d}\mathsf{S}(f)$ $=\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}\log f\Big).$
It remains to show the non-interaction conditions. We have, by the anti-symmetry of $\mathsf{J}$ and the definition of $\widetilde{\nabla}$:
$\mathsf{L}(f){\,\rm d}\mathsf{S}(f)$ $=-\nabla\cdot\big(f\mathsf{J}\nabla\log f\big)=0$
$\text{and}\quad\mathsf{M}(f){\,\rm d}\mathsf{E}(f)$ $=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}\frac{|v|^{2}}{2}\Big)=0.$
As mentioned in the introduction, by taking $x=x_{*}$, we formally obtain the GENERIC formulation for the classical inhomogeneous (kinetic) Landau equation. It is worth mentioning that the classical kinetic Boltzmann equation can also be formally cast into the GENERIC framework  [Ött97, Grm18].